### v1

# A Guide to Stochastic Optimisation for Large-Scale Inverse Problems

Abstract
Stochastic optimisation algorithms are the de facto standard for machine learning with large amounts of data. Handling only a subset of available data in each optimisation step dramatically reduces the per-iteration computational costs, while still ensuring significant progress towards the solution. Driven by the need to solve large-scale optimisation problems as efficiently as possible, the last decade has witnessed an explosion of research in this area. Leveraging the parallels between machine learning and inverse problems has allowed harnessing the power of this research wave for solving inverse problems. In this survey, we provide a comprehensive account of the state-of-the-art in stochastic optimisation from the viewpoint of inverse problems. We present algorithms with diverse modalities of problem randomisation and discuss the roles of variance reduction, acceleration, higher-order methods, and other algorithmic modifications, and compare theoretical results with practical behaviour. We focus on the potential and the challenges for stochastic optimisation that are unique to inverse imaging problems and are not commonly encountered in machine learning. We conclude the survey with illustrative examples from imaging problems to examine the advantages and disadvantages that this new generation of algorithms bring to the field of inverse problems.
: Inverse Problems
  • June 10, 2024
Keywords: Inverse Problems, Imaging, Variational Regularisation, First-Order Algorithms, Large-Scale Optimisation, Stochastic Optimisation
1   Inverse Problems Meets Optimisation
This review is concerned with the efficient solution of optimisation problems in the context of inverse problems. This is particularly relevant when data sets are large or forward operators are costly to evaluate. In order to set the scene, we are interested in recovering an object $u^{\dagger}\in\mathcal{U}$ from indirect noisy measurements $v\in\mathcal{V}$. We assume that the relationship between $u$ and $v$ can be approximated by the forward model
$K(u^{\dagger})\approx v,$ (1)
where $K:\mathcal{U}\to\mathcal{V}$ is a mapping between Hilbert spaces [^1] While some mathematical modelling for inverse problems is formulated on Banach spaces, see e.g. [BB18], for optimisation, it is vastly more convenient to consider Hilbert spaces [BC11, HH14]. The dimension of such spaces is not as important and many results are dimension independent. $\mathcal{U}$ and $\mathcal{V}$. Many tasks can be described through this expression, such as deblurring in image processing, tomographic reconstruction in imaging or regression in statistics, to name a few. Concrete applications in tomography include computed tomography (CT) [HJL21], positron emission tomography (PET) [OF97] and magnetic resonance imaging (MRI) [Fes10].
In this review we primarily consider linear inverse problems, where the mapping $K$ is a linear operator and we denoting its action on elements of $\mathcal{U}$ by $Ku$. Even though nonlinearity introduces additional mathematical and computational challenges, many techniques presented in this paper can be readily extended to nonlinear problems.
In applications both the solution $u$ and data $v$ are typically discretised, and we seek to recover an element in $\mathbb{R}^{d}$ from measurements in $\mathbb{R}^{s}$. Complex numbers can be readily incorporated by identifying $\mathbb{C}$ with $\mathbb{R}^{2}$. An entry $v_{i}$ represents the $i$-th measurement, e.g. of a digital detector, and $u_{i}$ corresponds to the $i$-th parameter of the solution, e.g. the value of a particular image voxel.
Many modern applications have in common that either $d$ and $s$ are both large, or that the forward operator $K$ is expensive to evaluate. Consequently, computing an approximate solution to the inverse problem (1) is computationally challenging. These two issues are generally highly related as e.g. the cost of matrix-vector products scales linearly with both the number of input and output dimensions. To illustrate the point of high dimensions, consider the PET-MR scanner Siemens Biograph mMR where the uncompressed data (so called “span 1”) has more than 350 million elements and images are commonly modelled with around 20 million voxels. It is important to highlight, that the words “large” and “expensive” are relative, and have to be compared to the available resources in a specific application.
Instead of directly solving (1), it is common to approximate the solution via an optimisation problem. The typical choice is
$\min_{u\in\mathcal{U}}D(Ku,v),$ (2)
where $D:\mathcal{V}\times\mathcal{V}\to\mathbb{R}_{\infty}:=\mathbb{R}\cup\{\infty\}$ measures how closely $Ku$ approximates $v$, and is commonly called the data-fidelity or data fit term. The data-fidelity term $D$ is typically a non-negative function minimised by identical arguments, $\inf_{v^{\prime}}D(v^{\prime},v)=D(v,v)=0$. Allowing $D$ to take on the value $\infty$ is sometimes needed to extend the definition of $D$ to a vector space (e.g. see the Kullback–Leibler divergence below) or to model constraints. See [BL18] for a discussion on extended real numbers $\mathbb{R}_{\infty}$.
An alternative viewpoint is to model, and then maximise, the likelihood of observing $v$ given $Ku$. The data-fidelity can then be defined as the negative log-likelihood and the sought estimator is the one maximising the likelihood, see e.g. [BB98, BB18] for more dedicated discussions.
Most commonly used data fidelities are separable. In the finite-dimensional data setting, $\mathcal{V}=\mathbb{R}^{s}$, this means that they can be written as
$D(v^{\prime},v)=\sum_{i=1}^{s}d(v_{i}^{\prime},v_{i}).$ (3)
The function $d:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{\infty}$ can be determined using statistical properties of the data or the measurement noise. For example, using the Gaussian, Laplace and Poisson distributions leads to least squares, absolute differences and Kullback–Leibler divergence, respectively, defined as follows
  • Least squares
    $d(v_{i}^{\prime},v_{i})=\frac{1}{2\sigma_{i}^{2}}(v_{i}^{\prime}-v_{i})^{2}.$
  • Absolute differences
    $d(v_{i}^{\prime},v_{i})=|v_{i}^{\prime}-v_{i}|.$
  • Kullback–Leibler divergence
    $d(v_{i}^{\prime},v_{i})=\operatorname{KL}(v_{i}|v_{i}^{\prime})=\begin{cases}v_{i}^{\prime}-v_{i}+v_{i}\log(v_{i}/v_{i}^{\prime})&\text{if $v_{i}^{\prime}>0$ and $v_{i}>0$},\\ v_{i}^{\prime}&\text{if $v_{i}^{\prime}>0$ and $v_{i}=0$},\\ \infty&\text{else}.\end{cases}$
While the Gaussian assumption is most frequently applied, Poisson distributed data has also gained a lot of attention, see e.g. [SST10, Bar00, HW16].
1.1   Variational Regularisation
The original inverse problem (1) is ill-posed, in the sense that it can have no solutions, infinitely many solutions or even if the solution exists and is unique, it is highly sensitive to the noise in the measurements. Solving the optimisation problem (2), which minimises the fidelity of the reconstruction to the measurements, instead of (1), can help overcome the nonexistence of the solution. However, it inherits other issues, such as non-uniqueness of solutions and noise amplification. For example, if $K$ is ill-conditioned, small changes in the observations (for example due to noise) can lead to large changes in the reconstructed solution. Moreover, only minimising the data fidelity will result in reconstructions that fit to noise as well. A common approach to address this is to add a regulariser $R:\mathcal{U}\to\mathbb{R}_{\infty}$ to the optimisation objective and instead solve
$\min_{u\in\mathcal{U}}\left\{D(Ku,v)+\lambda R(u)\right\},$ (4)
where $\lambda>0$ is called the regularisation parameter. This concept is known as variational regularisation and has been extensively studied extensively over the past decades. Many regularisers $R$ have been used over the years, see [SGG+08, BB18] and the references therein. Instead of providing a comprehensive overview, we will discuss representative regularisers with properties relevant for inverse problems.
If $\mathcal{U}=\mathbb{R}^{d}$, the simplest regulariser is the squared 2-norm: $R(u)=\frac{1}{2}\|u\|^{2}_{2}=\frac{1}{2}\sum_{i}u_{i}^{2}$. This is often referred to as Tikhonov or Miller regularisation [IJ14, Mil70] and can be used to stabilise the reconstruction and make the modelled solution unique. Alternatively, using the 1-norm, $R(u)=\|u\|_{1}=\sum_{i}|u_{i}|$, promotes a specific property in the solution: sparsity, i.e. only few coefficients of the solution are nonzero.
Regularising with 1- and squared 2-norm does not capture well the structure of desired solutions generally encountered in imaging applications. They promote reconstructions with small entries which is not desirable for images. Moreover, it does not take into account local neighbourhood or smoothness features within the image. Therefore, regularisers that penalise differences between neighbouring pixels are better suited for imaging tasks. This can be achieved by not penalising $u$ directly but rather its gradient: $R(u)=\frac{1}{2}\|\nabla u\|^{2}_{2}$. However, this is in strict terms well-defined only for smooth functions $u$. An alternative perspective is to interpret $\nabla$ as a linear operator that computes (spatial) finite differences of a discretised vector $u$. We do not to want delve deeper into the details, but instead refer to e.g. [BO13] or [CP16] for more details in the continuous or discrete setting, respectively.
Using the squared $2$-norm promotes similarity between neighbouring pixels and can cause blurring on the edges of an object. In order to preserve the edges it is common to replace the squared $2$-norm of the gradient with the $1$-norm, as
$R(u)={|\kern-1.125pt|}\nabla u{|\kern-1.125pt|}_{1}.$ (5)
This defines the total variation (TV), which is a particularly good model if we believe that the desired solution can be well described as a piece-wise constant function, a common model for natural images.
It is known that TV introduces staircasing artefacts when an image has piece-wise linear structures, see e.g. [CCN07]. To circumvent this problem, several high-order variants of TV have been proposed, see e.g. [CL97, BKP10]. Perhaps the most popular choice is the total generalised variation (TGV) [BKP10], defined as
$R(u)=\inf_{w}\left\{{|\kern-1.125pt|}\nabla u-w{|\kern-1.125pt|}_{1}+\beta{|\kern-1.125pt|}\nabla w{|\kern-1.125pt|}_{1}\right\},$ (6)
for some $\beta>0$. Here $w$ is the vector field of the piece-wise linear component of $u$, while $\nabla u-w$ is the vector field of the piece-wise constant component of $u$.
An alternative to the non-smooth TV is the so called Huber TV, see e.g. [CP16], which replaces the $1$-norm with a smooth approximation. In the discretised setting it can be written as
$R(u)=\sum_{i}h_{\gamma}(\|(\nabla u)_{i}\|_{2}),$
where the sum is over all components of the gradient field $\nabla u$ and
$h_{\gamma}(t)=\begin{cases}t&\text{if }|t|>\gamma,\\ \frac{1}{2\gamma}t^{2}+\frac{\gamma}{2}&\text{if }|t|\leq\gamma.\end{cases}$
Huber TV is differentiable, with a $\|\nabla\|^{2}/\gamma$-Lipschitz continuous gradient, and in several applications it overcomes the staircasing artefact that TV introduces.
Another source of regularisation can be formulated as constraints. In contrast to the previously discussed models, which encourage the solution to have certain properties, constraints can be used to instead enforce the desired properties. The most popular constraints in imaging are box constraints, where we only seek solutions within a $d$-dimensional box
$C=\{x\in\mathbb{R}^{d}:\ell_{i}\leq x_{i}\leq r_{i},i=1,2,\ldots,d\}.$
The most relevant constraints of this type are the non-negativity constraints, where one takes $\ell_{i}=0$ and $r_{i}=\infty$.
Remark 1 (Statistical Perspective)
Variational regularisation (4) can be articulated from a Bayesian perspective in terms of maximum a posteriori (MAP) estimation [Stu10]. Given a density $p(u)$ for the probabilistic model of $u$, MAP considers maximising the posterior probability $p(u|v)\propto p(v|u)p(u)$, which is equivalent to minimising its negative logarithm, i.e.
$\min_{u}\left\{-\log p(v|u)-\log p(u)\right\}.$
For example, if we assume that the data is Gaussian distributed with mean $Ku$ and covariance $\sigma^{2}I$ and that the solution is Gaussian distributed with zero mean and covariance $\lambda^{-1}I$, the MAP can be written as $\operatornamewithlimits{argmin}_{u}\left\{\sigma^{-2}\|Ku-v\|^{2}+\lambda\|u\|^{2}\right\}$.
The separability of the data fit (3) can be motivated from a statistical perspective, and follows for instance from assuming that the data points are conditionally independent given the estimated solution.
1.2   Optimisation Template
As we just discussed, inverse problems can be solved via optimisation in the context of variational regularisation (4). All aforementioned optimisation problems are special cases of the optimisation template
$\min_{x\in\mathcal{X}}\Bigl{\{}\Phi(x)=f(Ax)+g(x)+h(x)\Bigr{\}}.$ (7)
Here the functions $f:\mathcal{Y}\to\mathbb{R}_{\infty}$ and $g:\mathcal{X}\to\mathbb{R}_{\infty}$ are convex and prox-friendly, i.e. they have the property that their proximal operators, see (12) for a definition, can be either evaluated in closed-form or efficiently approximated. The function $h:\mathcal{X}\to\mathbb{R}$ is convex and $L$-smooth, i.e. it is continuously differentiable and its gradient has Lipschitz constant $L>0$, see (10). Lastly, $A:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator. Note that we do not want to (or can) compute the proximal operator of $f\circ A$, but instead want to decouple the function $f$ from the operator $A$, and treat them separately. Techniques to do so are discussed in Section 2.
The conversion from a given minimisation problem for variational regularisation to the optimisation template is not unique as we show in the following example.
Example 1.1
We consider computing a TV-regularised solution as
$\min_{u\in\mathcal{U}}\left\{\frac{1}{2}\|Ku-v\|_{2}^{2}+\lambda\|\nabla u\|_{1}\right\}.$
This can be written via our template (7) with $x=u$, $\mathcal{X}=\mathcal{U}$, and
  • (i)
    $f(y)=\lambda\|y\|_{1},A=\nabla,g(x)=0,h(x)=\frac{1}{2}\|Kx-v\|_{2}^{2}$;
  • (ii)
    $f(y)=0,A=0,g(x)=\lambda\|\nabla x\|_{1},h(x)=\frac{1}{2}\|Kx-v\|_{2}^{2}$;
  • (iii)
    $f(y)=0,A=0,g(x)=\|\nabla x\|_{1},h(x)=\frac{1}{2\lambda}\|Kx-v\|_{2}^{2}$.
The three formulations above (and many more) are all mathematically valid, and used in many applications but lead to algorithms with very different properties. For most of these choices the proximal operators of $f$ and $g$ can be computed in closed-form. For (ii) and (iii), it is common to compute the proximal operator of the total variation $\|\nabla x\|_{1}$ via warm-started fast gradient projection (FGP) iterations [BT09], see e.g. [EB16] for an MRI application.
The next example illustrates that the optimisation variable $x$ does not need to coincide with the inverse problems variable $u$.
Example 1.2
We consider the same inverse problem as in Example 1.1, but replace the regulariser with TGV (6), i.e. we aim to solve
$\min_{u,w}\left\{\frac{1}{2}\|Ku-v\|_{2}^{2}+\lambda{|\kern-1.125pt|}\nabla u-w{|\kern-1.125pt|}_{1}+\lambda\beta{|\kern-1.125pt|}\nabla w{|\kern-1.125pt|}_{1}\right\}.$
The minimiser can then be computed via template (7) with $x=(u,w)$ and
$f(y_{1},y_{2})=\lambda\|y_{1}\|_{1}+\lambda\beta\|y_{2}\|_{1},\quad A=\begin{pmatrix}\nabla&-I\\ 0&\nabla\end{pmatrix},\quad g(x)=0,\quad h(u,w)=\frac{1}{2}\|Ku-v\|_{2}^{2}.$
Most notably, the optimisation variable $x$ consists of the image to be reconstructed $u$ and the vector-field $w$ in the definition of TGV.
For the third example, we consider how can constraints be incorporated into the optimisation template.
Example 1.3
We consider the same problem as in Example 1.1, but this time we add the non-negativity constraint
$\min_{u\geq 0}\left\{\frac{1}{2}\|Ku-v\|_{2}^{2}+\lambda\|\nabla u\|_{1}\right\}.$
This can be set into our template (7) with $u=x$, $f(y)=\|y\|_{1},A=\nabla,h(x)=\frac{1}{2}\|Kx-v\|_{2}^{2}$ and $g(x)=\iota_{[0,\infty)^{d}}(x)$ is the indicator function of the non-negative orthant (see (11) for the definition). Since all $x$ that violate the constraint take the value $\infty$, this choice for $g$ guarantees that the optimal solution is indeed non-negative.
Though the above examples are mostly concerned TV regularisation, they are readily extended to other types of regularisation such as Huber TV, or Wavelet regularisation (by replacing the gradient operator $\nabla$ with a wavelet transform $W$).
1.3   Iterative Regularisation
Variational regularisation (4) gives a mathematically modelled solution, not an algorithm. An often used algorithm to solve (7) when $f=0$ is
$x^{(k+1)}=\operatorname{prox}_{\tau_{k}g}\left(x^{(k)}-\tau_{k}\nabla h(x^{(k)})\right).$ (8)
This algorithm is a fundamental building block in optimisation. It has many names, including proximal gradient descent and forward-backward splitting. If $g=0$ then (8) reduces to the classical gradient descent. We will discuss this and other algorithms to solve (7) in more detail in Section 2. For well chosen step-sizes, iterations (8) converge to a solution defined via variational regularisation.
An alternative is to not include a regulariser directly and instead stop the iterations early. For example, the Landweber iterations
$u^{(k+1)}=u^{(k)}-\tau K^{*}(Ku^{(k)}-v),$ (9)
with $u^{(0)}=0$ and well-chosen $\tau$ and stopping index $k$, converge to the minimal-norm solution of (1) if it exists [Lan51]. Choosing a good stopping criteria is an important and task dependent problem. Landweber iterations for a nonlinear forward operator $K$ were analysed in [HNS95]. Turning an optimisation method into a regularisation method is commonly referred to as iterative regularisation, which has been used for linearised Bregman iterations [YOGD08, Yin10] and their non-convex variants [BBES21], dual algorithms [GRV18], primal-dual algorithms [MMRV21], and analysed in Banach spaces [SLS06], to name a few.
1.4   Further Reading
Advanced Modelling
The present survey focuses on the most commonly used mathematical models for variational regularisation. However, this only scratches the surface of what can be done in this context. For example, a drawback of variational regularisation is the bias introduced by the regulariser, particularly when the data is not very informative. The bias can be reduced by the two-step technique in [BBRS17] or iterative regularisation [BB18]. Another interesting direction is the coupling of multiple measurements, which can correspond to different modalities or energy channels [Ehr23, AET21] or by considering the problem as a dynamic inverse problem over time [SHB18, HÖS21]. Such forward models require tailored mathematical modelling that can be done in the context of variational regularisation.
Machine Learning
Recent years have seen a rise of learned regularisers for inverse problems [AMOS19], for example dictionary learning [TF11], fields-of-expert model [RB05], gradient-step denoisers [HLP21, TMTS24b], regularisation-by-denoising [REM17], deep image prior with TV [BLS20, BAL+24], generative regularisers [DCE24], approaches based on the input-convex neural networks (ICNN) [AXK17] or network Tikhonov (NETT) [LSAH20]. Among them, a focus is on learning informative yet convex regularisation, which utilise specially parameterised neural networks, such as the ICNN [MHÖ+23] and convex ridge regularisers [GNB+22].
1.5   Other Reviews
To the best of our knowledge this is the first review that focuses on stochastic optimisation algorithms for inverse problems. However, there have been other reviews concerned with the efficient computation of inverse problems, some of which also focus on optimisation approaches. In [CFKS13] and [CKN15], the authors discuss the efficient solution of large-scale inverse problems focusing on deterministic algorithms from numerical linear algebra. The reviews [BSS16],[CP16], and [Val16] are concerned with efficient first-order optimisation algorithms for inverse problems and imaging. While we also discuss first-order algorithms in detail, our particular focus is on stochastic approaches to overcome the large-scale nature of the problem. The paper [BBM23] surveys stochastic methods for training neural networks for image classification but does not consider inverse problems. The connection of convex optimisation to machine learning and AI for medical imaging is discussed in [XN22]. They do touch upon stochastic optimisation but do not go into much detail or specifics that are of interest for inverse problems.
1.6   Mathematical Preliminaries and Notation
[…]
By $(a,b]$ we denote the interval $a<x\leq b$. We write $x_{i}$ for the $i$-th entry of a vector $x\in\mathbb{R}^{d}$. For $1\leq p<\infty$ we denote the $p$-norm of a vector $x\in\mathbb{R}^{d}$ by $\|x\|_{p}=(\sum_{i=1}^{d}|x_{i}|^{p})^{1/p}$.
A function $f$ is called convex if
$f(\alpha x_{1}+(1-\alpha)x_{2})\leq\alpha f(x_{1})+(1-\alpha)f(x_{2}),~{}\forall\alpha\in[0,1]$
and it is strictly convex if the above equality is strict. It is $\mu$-strongly convex if
$f(x)-\frac{\mu}{2}\|x\|^{2}$
is convex. It is $L$-Lipschitz continuous if for some $L\geq 0$ it holds that
$\|f(x_{1})-f(x_{2})\|\leq L\|x_{1}-x_{2}\|.$ (10)
A function $f$ is called $L$-smooth if it is differentiable and has $L$-Lipschitz continuous gradient. If $f$ is twice differentiable and $\mu$-strongly convex, then all the eigenvalues of its Hessian are all larger than $\mu$ (i.e. $\nabla^{2}f(x)\succeq\mu I)$, and it is $L$-smooth if they are all bounded by $L$ (i.e. $\nabla^{2}f(x)\preceq LI$).
A function $f$ is called proper if its effective domain $\{x\in\mathcal{X}:f(x)<\infty\}$ is nonempty. It is also called closed, if its sublevel set $\{x\in\mathcal{X}:f(x)\leq\gamma\}$ is closed for each $\gamma\in\mathbb{R}$.
When $f$ is proper closed and convex, its subdifferential is defined by
$\partial f(x)=\{p\in\mathcal{X}:f(z)\geq f(x)+\langle p,z-x\rangle,\text{ for all }z\in\mathcal{X}\}.$
The Fenchel (or convex) conjugate of a function $f$ is defined as
$f^{\ast}(y)=\sup_{x\in\mathcal{X}}\left\{\langle x,y\rangle-f(x)\right\}.$
Note that $f^{\ast}$ is closed convex regardless the convexity of $f$. The indicator function $\iota_{\mathcal{C}}$ of a set $\mathcal{C}\subset\mathcal{X}$ is defined as
$\iota_{\mathcal{C}}(x)=\begin{cases}0,&\text{ if }x\in\mathcal{C},\\ \infty,&\text{otherwise}.\end{cases}$ (11)
The proximal operator with parameter $\tau>0$ of a proper closed convex function $f:\mathcal{X}\rightarrow\mathbb{R}_{\infty}$ is the mapping $\operatorname{prox}_{\tau f}:\mathcal{X}\rightarrow\mathcal{X}$ defined by
$\operatorname{prox}_{\tau f}(z)=\operatornamewithlimits{argmin}_{x\in\mathcal{X}}\left\{f(x)+\frac{1}{2\tau}\|x-z\|^{2}\right\}.$ (12)
Due to the convexity of $f$ (and thus the strong convexity of the objective), the minimiser always exists and is unique. If $\mathcal{X}=\mathbb{R}^{d}$ with the Euclidean norm, then the proximal operator of many convex and non-convex functions, e.g. norms and indicator functions, has a closed-form expression [CP11b], and is known for many other functions [CCCP]. For instance, the proximal operator of the indicator function of the non-negative orthant $[0,\infty)^{d}$ has a particularly simple expression $\left[\operatorname{prox}_{\tau\iota_{[0,\infty)^{d}}}(x)\right]_{i}=\max(x_{i},0)$.
2   Deterministic Algorithms for Variational Regularisation
Developing efficient solvers to tackle variational optimisation problems has been an incredibly active and productive research area, especially since the new millennium. Among the existing methodologies, first-order methods, which utilise only the first-order differential information of the objective function, remain the most popular.
In this section we provide an overview of first-order methods, from foundational principles to recent algorithmic advancements. In Section 2.1 we introduce the basic building blocks and design principles for first-order methods. In Section 2.2 we present representative first-order algorithms for minimising variants of (7). We conclude in Section 2.3 with a discussion of further topics: higher-order schemes, non-Euclidean methods, nonconvex optimisation and nonlinear problems
We emphasise that this is a rich and rapidly evolving field of research, and consequently, our presentation is not comprehensive. We refer to monographs [BC11, Bec17] and reviews [CKCH23, BB09, BB18, BBES21] for more details.
2.1   Building Blocks of First-Order Algorithms
Recall the optimisation template from Section 1.2
$\min_{x\in\mathcal{X}}\Big{\{}\Phi(x)=f(Ax)+g(x)+h(x)\Big{\}},$
where functions $f$ and $g$ are proper closed convex and moreover prox-friendly, $A$ is a linear operator and $h$ is a convex and $L$-smooth function. As discussed in the introduction, the above template covers a wide range of optimisation problems from inverse imaging problems.
The difficulties of solving (7) come from several sources, such as nonsmoothness, function-operator composition, and the summation of several terms. Divide and conquer is a well-established and often used principle for designing algorithms to address these difficulties. It consists of decomposing the original problem into simpler sub-problems, tackling each sub-problem separately, and assembling all the components into a provable and implementable numerical scheme.
2.1.1   Basic Elements
The divide and conquer principle relies on having a range of approaches for addressing individual sub-problems that may arise. In terms of the optimisation problem (7), the crux of the matter is managing $h$, $g$ and $f\circ A$.
Gradient Descent
Gradient descent (GD) is one of the oldest iterative methods for unconstrained optimisation, dating back to at least the middle of the 19th century. When minimising a differentiable objective $h$, GD iteratively performs a step in the direction of the negative gradient at the current iterate
$x^{(k+1)}=x^{(k)}-\tau_{k}\nabla h(x^{(k)}),$ (13)
where $\tau_{k}>0$ is the step-size. As the name suggests, GD is a descent method that reduces function values in each iteration $h(x^{(k+1)})\leq h(x^{(k)})$ assuming that the step-sizes $\tau_{k}$ are well-chosen [^2] The optimisation literature often focuses on linear convergence rate for the strongly convex case. Here, we put more emphasis on the merely convex case, since in practice the optimisation problems associated to inverse problems are often only strongly convex with a very small convexity parameter.
Proximal Point Algorithm
When minimising the non-smooth function $g$, one can consider GD by replacing the gradient with a sub-gradient. However, the resulting sub-gradient method is no longer a descent algorithm for constant step-sizes [Bec17]. A much more popular approach is to use the proximal operator of $g$. For a proper, closed and convex function $g$ the corresponding proximal point algorithm (PPA) [Roc76], or the proximal minimisation algorithm, takes the form
$x^{(k+1)}=\mathrm{prox}_{\tau_{k}g}(x^{(k)}):=\operatornamewithlimits{argmin}_{x\in\mathcal{X}}\left\{\tau_{k}g(x)+\frac{1}{2}{|\kern-1.125pt|}x-x^{(k)}{|\kern-1.125pt|}^{2}\right\},$ (14)
where $\tau_{k}>0$ is the step-size. Convergence of PPA to a minimiser is ensured provided the step-size does not vanish too quickly, $\sum_{k}\tau_{k}=\infty$. Note that each iteration requires solving an optimisation problem to compute the proximal operator. However, the proximal operator of many functions is known in closed-form [CP11b, CCCP].
Remark 2
GD and PPA can be interpreted as different discretisations of the gradient flow $x^{\prime}(t)=-\nabla g(x(t)),t\geq 0$. Applying the explicit Euler scheme to the gradient flow yields GD (13): $(x^{(k+1)}-x^{(k)})/\tau=-\nabla g(x^{(k)})$. On the other hand, applying the implicit Euler scheme $(x^{(k+1)}-x^{(k)})/\tau=-\nabla g(x^{(k+1)})$, requires solving an equation $x^{(k+1)}+\tau\nabla g(x^{(k+1)})=x^{(k)}$, to compute $x^{(k+1)}$. This in turn is equivalent to $x^{(k+1)}=\mathrm{prox}_{\tau g}(x^{(k)})$, which is PPA (14).
Acceleration
Acceleration is an essential component of any effective first-order method. The most successful acceleration technique is “inertial acceleration” or “momentum”, dating back to the heavy-ball method [Pol64] which can significantly speed up GD, but lacks global acceleration guarantees. Nesterov’s accelerated gradient (NAG) method was the first inertial acceleration that provably improves the convergence rate of GD for smooth convex objectives [Nes83]. Fast iterative soft-thresholding algorithm (FISTA) extends this speed up to nonsmooth convex objectives [BT09]. We refer to [ABC23] for a more comprehensive treatment of the development of inertial acceleration over the past decade. Other acceleration techniques include preconditioning [PC11, LXY21], vector extrapolation [And65, SdB16] and so on. We refer to the monograph [dST21] on acceleration for further details.
Remark 3
Similar to GD, NAG also has an ODE interpretation. In [SBC16], the authors showed that the limit of NAG iterations is a second-order differential equation: $x^{\prime\prime}(t)+({3}/{t})x^{\prime}(t)=-\nabla g(x(t))$.
Function-Operator Decomposition
Now we turn to the composition term $f\circ A$, whose proximal operator in general is difficult to compute. A notable exception is when $AA^{\ast}=\alpha I,~{}\alpha>0$, in which case the proximal operator of $f\circ A$ can be computed as $\operatorname{prox}_{f\circ A}(x)=x+\tfrac{1}{\alpha}A^{\ast}\big{(}\operatorname{prox}_{\alpha f}(Ax)-Ax\big{)}$ [Bec17]. For the general case this issue can be circumvented by one of two popular approaches: duality or method of multiplier. The former uses the Fenchel–Moreau theorem, which states that the bi-conjugate of a proper, closed convex function is equal to itself. This gives
$f(Ax)=f^{\ast\ast}(Ax):=\sup_{y\in\mathcal{Y}}\big{\{}\langle Ax,y\rangle-f^{\ast}(y)\big{\}},$ (15)
and thus the optimisation problem can be recast as a saddle-point problem, and then solved by the primal-dual algorithm PDHG (see Algorithm 3). An alternative is to decompose $f\circ A$ via an auxiliary variable into a constrained problem
$\min_{x\in\mathcal{X},\ z\in\mathcal{Z}}\big{\{}f(z)\quad\text{ such that }\quad z=Ax\big{\}}.$
By adding the constraint back to the objective via a Lagrange multiplier, the problem can be solved by dual ascent, method of multipliers [BPC+11], or the alternating direction method of multipliers (ADMM) (see Algorithm 4). PDHG and ADMM are deeply connected, and equivalent to each under some conditions, see [CP11a, CP16, OV20] for discussions.
Coordinate Descent
Instead of updating all the entries of the optimisation variable $x$ in each iteration, coordinate descent (CD) methods update only one or a small subset of coordinates:
$x^{(k+1)}=x^{(k)}-\tau_{k}\nabla_{B_{k}}h(x^{(k)})e_{B_{k}}=x^{(k)}-\tau_{k}\sum_{i\in B_{k}}\nabla_{i}h(x^{(k)})e_{i}.$ (16)
Here $B_{k}$ is a subset of the coordinates of $x$, $e_{i}$ is the $i$-th canonical basis vector and $\nabla_{i}h(x^{(k)})=\frac{\partial h(x^{(k)})}{\partial x_{i}}$ denotes the directional derivative along $i$-th coordinate. When $B_{k}$ is a singleton, (16) is the basic CD scheme, otherwise it is called block coordinate descent. CD-based methods are commonly used for problems that have a separable structure or that are coordinate-friendly [STXY16]. Popular deterministic choices for the order in which the coordinates are updated include cyclic and fixed permutation orders. Randomised orders are discussed in Section 3.3. The introduction of CD to optimisation is due to [BT15] [^3] The first edition of the book was published in 1989.. Convergence results were first shown for differentiable objectives [LT92], and later extended to non-differentiable objectives [Tse01]. We refer to [Nes12, BT13, RT14] for theoretical advances of CD algorithms.
2.1.2   Designing a First-Order Algorithm
With the above basic elements at hand, designing a first-order methods consists of the following steps:
  • Problem decomposition The original optimisation problem is decomposed into simpler sub-problems, according to properties of the individual functions.
  • Sub-problem processing and assembly Each sub-problem is assigned an approach to tackle it (using e.g. Table 1 as a guideline). The resulting components are then assembled to form an iterative scheme. In principle, optimality conditions can serve as a reference for the assembly of the algorithm.
  • Refinement The iterative scheme might require further improvements to address an existing issue. First, one or several of its components might require solving an additional optimisation problem, e.g. for the computation of the proximal operator. Second, theoretical convergences guarantees might be either lacking or too restrictive, and thus the approach ought to be modified. Third, theoretical or practical convergence may be slow, and thus require acceleration or other techniques to improve the performance.
Property Method References
Smoothness $x^{(k+1)}=x^{(k)}-\tau_{k}\nabla h(x^{(k)})$ [Kan60, BC11]
Prox-friendly $x^{(k+1)}=\mathrm{prox}_{\tau_{k}g}(x^{(k)})$ [Mor65, PB14]
Composition Duality: $f(x)=\sup_{y}\left\{\langle y,Ax\rangle-f^{\ast}(y)\right\}$ [Cha04, CP11a]
Constraint: $\min_{x,z}\big{\{}f(z)~{}~{}~{}\textrm{s.t.}~{}~{}z=Ax\big{\}}$ [GM76, BPC+11]
Coordinate friendly $x^{(k+1)}=x^{(k)}-\tau_{k}\nabla_{i_{k}}h(x^{(k)})e_{i_{k}}$ [BT15, Wri15]
Table 1: Foundational elements in designing first-order algorithms
It is important to emphasise that the above steps are not necessarily unique. In the following we present classical first-order algorithms that can be applied to solve instances of (7). We start with simpler cases where only two of the terms in (7) are nontrivial, and then move to the more general three-term case.
2.2   Advanced First-Order Algorithms
In this section we present representative first-order algorithms for minimising two-term and three-term optimisation problems.
2.2.1   Approaches for a Two-Term Sum
Minimisation of a two-term sum of functions is widely studied with several popular approaches in the literature. Moreover, approaches for two-term problems often serve as building blocks for more advanced methods.
Proximal Gradient Descent
When $f$ vanishes, (7) reduces to a two-term sum
$\min_{x\in\mathcal{X}}\left\{\Phi(x)=g(x)+h(x)\right\},$ (17)
where $h$ is smooth and $g$ is prox-friendly. The most popular scheme to solve (17) is the proximal gradient descent (PGD), see Algorithm 1, also known as the forward-backward splitting method [CW05]. PGD performs a gradient descent step on the smooth term $h$, and then evaluates the proximal operator of $g$.
Input: $x^{(0)}\in\mathcal{X}$, $\tau>0$.
1 for $k=0,1,...$ do
2       $x^{(k+1)}=\mathrm{prox}_{\tau g}\big{(}x^{(k)}-\tau\nabla h(x^{(k)})\big{)}$
3 end for
Algorithm 1 Proximal Gradient Descent
PGD is the simplest extension of GD to the nonsmooth case, and it shares similar convergence properties: provided that $h$ is $L$-smooth and the step-size satisfies $\tau<2/L$, PGD converges at a rate $\mathcal{O}(1/k)$ in the objective [BT09] [^4] The optimisation literature often considers the strongly convex case, too. We omit it here since in practice the optimisation problems associated to inverse problems are merely strongly convex.
Accelerated PGD
In spite of its simplicity, PGD is in practice hindered by its slow $\mathcal{O}(1/k)$ convergence rate. Fast iterative soft-shrinkage algorithm (FISTA) [BT09], extends the acceleration methodology of NAG to the non-smooth convex case.
Input: $x^{(0)}=x^{(-1)}\in\mathcal{X}$, $\tau>0$, $t_{0}=1$.
1 for $k=0,1,...$ do
2       $t_{k+1}=\frac{1+\sqrt{1+4t_{k}^{2}}}{2},~{}a_{k}=\frac{t_{k}-1}{t_{k+1}}$
3      $y^{(k+1)}=x^{(k)}+a_{k}(x^{(k)}-x^{(k-1)})$
4      $x^{(k+1)}=\mathrm{prox}_{\tau g}\big{(}y^{(k+1)}-\tau\nabla h(y^{(k+1)})\big{)}$
5 end for
Algorithm 2 FISTA
Provided that $h$ is convex and $L$-smooth and $\tau\leq 1/L$, FISTA improves the convergence rate in the objective value to $\mathcal{O}(1/k^{2})$. It can be shown that the iterates of FISTA converge to a global minimiser with a modified update rule $a_{k}=(k-1)/(k+d),d>3$ [CD15]. Moreover, the authors also show that the objective function convergence speed is indeed $o(1/k^{2})$.
Despite the theoretical advantage, in practice FISTA is not necessarily faster than PGD, especially when high precision solution is sought. The reason is that FISTA is not a descent method and shows (periodic) oscillatory behaviour, typically when the problem is (locally) strongly convex. To circumvent this problem, restarting schemes can significantly improve the performance of FISTA [OC15, SBC16].
Primal-Dual Splitting Methods
When $g$ vanishes, (7) reduces to
$\min_{x\in\mathcal{X}}\left\{\Phi(x)=f(Ax)+h(x)\right\},$ (18)
where $f$ and $g$ are both nonsmooth but prox-friendly. As previously discussed, the proximal operator of $f\circ A$ is in general not available in closed form and alternative approaches are needed to handle $f\circ A$.
Using the concept of duality, plugging (15) into (18) gives the saddle point problem
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{h(x)+\langle Ax,y\rangle-f^{\ast}(y)\right\}.$ (19)
Chambolle and Pock [CP11a] propose to solve (19) via an algorithm that alternates between a descent in the primal and an ascent in the dual on an extrapolated variable, see Algorithm 3 below for details.
Input: $x^{(0)}\in\mathcal{X},y^{(0)}\in\mathcal{Y}$, $\sigma,\tau>0$.
1 for $k=0,1,\ldots$ do
2       $x^{(k+1)}=\mathrm{prox}_{\tau g}\big{(}x^{(k)}-\tau A^{\ast}y^{(k)})$
3       $\bar{x}^{(k+1)}=2x^{(k+1)}-x^{(k)}$
4       $y^{(k+1)}=\mathrm{prox}_{\sigma f^{\ast}}\big{(}y^{(k)}+\sigma A\bar{x}^{(k+1)}\big{)}$
5 end for
Algorithm 3 Primal-Dual Hybrid Gradient Algorithm (PDHG)
This algorithm is known as the primal-dual hybrid gradient (PDHG) or the Chambolle–Pock algorithm. PDHG can be equivalently written as PPA on $\mathcal{X}\times\mathcal{Y}$ [HY12]. Hence, one can obtain the convergence of $(x^{(k)},y^{(k)})$ to a saddle point $(x^{\ast},y^{\ast})$ where $x^{\ast}$ is a solution of (17). The most known convergence result for PDHG is that the primal-dual gap converges as $\mathcal{O}(1/k)$ provided $\sigma\tau\|A\|^{2}<1$ [CP11a]. In a recent work [MCJH23] improved the step-size condition to $\sigma\tau\|A\|^{2}<4/3$. We also refer to [GLY+13, GLY15] for alternative adaptive step-size schemes.
Alternating Direction Method of Multipliers
Addressing $f\circ A$ through the method of multipliers converts (18) into a constrained problem
$\min_{x\in\mathcal{X},\ z\in\mathcal{Z}}\left\{\Phi(x,z):=f(z)+g(x)\right\}\quad\textrm{ such that}\quad Ax-z=0.$ (20)
The objective $\Phi$ is separable in $x$ and $z$, and the constraint enforces their coupling. The associated augmented Lagrangian reads
$\mathcal{L}(x,z;y)=f(z)+g(x)+\langle y,Ax-z\rangle+\frac{\tau}{2}\|{Ax-z}\|^{2}$ (21)
where $\tau>0$ and $w$ is the Lagrangian multiplier. To find a saddle-point of $\mathcal{L}(x,z;w)$, the Alternating Direction Method of Multipliers (ADMM, Algorithm 4) updates $z,x$, and $w$ one at a time.
Input: $x^{(0)}\in\mathcal{X},y^{(0)}\in\mathcal{Y},\tau>0$.
1 for $k=0,1,...$ do
2       $z^{(k+1)}\in\operatornamewithlimits{argmin}_{z\in\mathcal{Z}}~{}\mathcal{L}(x^{(k)},z,y^{(k)})$
3       $x^{(k+1)}\in\operatornamewithlimits{argmin}_{x\in\mathcal{X}}~{}\mathcal{L}(x,z^{(k+1)},y^{(k)})$
4       $y^{(k+1)}=y^{(k)}+\tau({Ax^{(k+1)}-z^{(k+1)}})$
5 end for
Algorithm 4 Alternating Direction Method of Multipliers (ADMM)
We refer to [GM75, GM76] for the early development of ADMM and [BPC+11] for a more modern survey. ADMM is equivalent to applying the Douglas–Rachford splitting method to the dual of (18), see [EB92, CP07].
The update for the auxiliary variable can be written as $z^{(k+1)}=\operatorname{prox}_{f/\tau}(Ax^{(k)}-z+\tfrac{1}{\tau}y^{(k)})$, and is thus uniquely determined. On the other hand, the minimisers of $\mathcal{L}(x,z^{(k+1)},y^{(k)})$, and thus $x^{(k+1)}$, need not to be unique, unless $A$ has full column-rank. The convergence analysis of ADMM is in general complicated. However, when $f$ and $h$ are closed and convex, and $\mathcal{L}$ admits a saddle point, convergence in the objective and the constraint can be shown [BPC+11]. The convergence of the primal sequences $x^{(k)}$ and $z^{(k)}$ is more difficult to establish but when $A$ is of full column-rank, the convergence of $(x^{(k)},z^{(k)})$ can be shown [EB92].
Remark 4
Instead of updating $z^{(k+1)}$ and $x^{(k+1)}$ sequentially, as in Algorithm 4, we can update them simultaneously
$(z^{(k+1)},x^{(k+1)})\in\operatornamewithlimits{argmin}_{x\in\mathcal{X},z\in\mathcal{Z}}~{}\mathcal{L}(x,z,y^{(k)}).$
The resulting method is called the augmented Lagrangian method [Bec17] or method of multipliers [BPC+11]. However, the simultaneous update has in general no closed-form expression. Hence, sequential (block) updates are preferred. This can be readily extended to problems with more than two blocks; see Section 2.2.2 for a discussion.
Remark 5
ADMM is traditionally derived to solve a variation of the problem (20) with a more general constraint: $Ax+Bz=u$. This requires only mild modifications of the Algorithm 4. However, the first step is no longer a proximal operator, but requires minimising $f(z)+\frac{\tau}{2}\|{Ax^{(k)}+Bz-u+\tfrac{1}{\tau}w^{(k)}}\|^{2}$, if $B^{\ast}B\neq I$. To solve this subproblem, there are two standard strategies. First strategy is to add a proximal term $\langle z-z^{(k)},Q(z-z^{(k)})\rangle$ with $Q=\frac{1}{\beta}I-B^{\ast}B$ where $\beta>0$ is such that $Q$ is symmetric positive definite. In doing so, the problem reduces to computing the proximal operator of $f$. The second strategy is to linearise the quadratic term at point $z^{(k)}$, which corresponds to applying one-step PGD to solve the corresponding subproblem.
2.2.2   Approaches for a Three-Term Sum
In previous section we discussed standard approaches for an objective decomposed into a sum of two terms. Most of the approaches presented therein can be extended to the three-term optimisation template (7). However, sometimes more flexibility is needed.
Condat–Vũ Algorithm
Condat and Vũ independently developed a primal-dual forward-backward algorithm for (7) as a generalisation of the PDHG, see Algorithm 5.
Input: $x^{(0)}\in\mathcal{X},y^{(0)}\in\mathcal{Y}$, $\tau,\sigma>0$
1 for $k=0,1,\ldots$ do
2       $x^{(k+1)}=\operatorname{prox}_{\tau g}\left(x^{(k)}-\tau\nabla h(x^{(k)})-\tau A^{*}y^{(k)}\right)$
3       $\overline{x}^{(k+1)}=2x^{(k+1)}-x^{(k)}$
4       $y^{(k+1)}=\operatorname{prox}_{\sigma f^{\ast}}\left(y^{(k)}+\sigma A\overline{x}^{(k+1)}\right)$
5 end for
Algorithm 5 Condat–Vũ Algorithm
The iterates of the Condat–Vũ algorithm converge if $\tau(\sigma\|A\|^{2}+L/2)<1$ where $L$ is the Lipschitz constant of the gradient of $h$ [Con13, Vũ13]. Note that each step requires one evaluation of $A,A^{*},\operatorname{prox}_{\tau g},\operatorname{prox}_{\sigma f^{\ast}}$ and the gradient $\nabla h$.
PD3O
The primal-dual three operator algorithm (PD3O) is another generalisation of PDHG and other two- and three-term algorithms. The difference between PD3O and Condat–Vũ is in the final term in the update for $y^{(k+1)}$, where PD3O adds the gradient momentum term $\tau(\nabla h(x^{(k)})-\nabla h(x^{(k+1)}))$.
Input: $x^{(0)}\in\mathcal{X},y^{(0)}\in\mathcal{Y}$, $\sigma,\tau>0$
1 for $k=0,1,\ldots$ do
2       $x^{(k+1)}=\operatorname{prox}_{\tau g}(x^{(k)}-\tau\nabla h(x^{(k)})-\tau A^{*}y^{(k)})$
3       $\overline{x}^{(k+1)}=2x^{(k+1)}-x^{(k)}+\tau(\nabla h(x^{(k)})-\nabla h(x^{(k+1)}))$
4       $y^{(k+1)}=\operatorname{prox}_{\sigma f^{\ast}}(y^{(k)}+\sigma A\overline{x}^{(k)})$
5
6 end for
Algorithm 6 Primal-Dual Three Operator Algorithm (PD3O)
PD3O requires one more evaluation of the gradient $\nabla h$ than Condat–Vũ, however, these costs can be amortised by storing $\nabla h(x^{(k+1)})$ and then using it in the next iteration.
It has been shown to converge for $\sigma\tau\|A\|^{2}<1$ and $\tau L<2$ [Yan18]. Importantly, the convergence condition decouples the conditions coming from PDHG and GD. This is in contrast to the Condat–Vũ algorithm where the convergence criteria couples the parameters $\tau$ and $\sigma$, with both $\|A\|^{2}$ and $L$.
Further Splitting Approaches
Methods presented in previous sections are only a small selection of first-order schemes for solving various instances of the optimisation template (7) but listing all the existing algorithms is beyond the scope of this paper. Other important algorithms for the two-term sum are the Douglas–Rachford splitting algorithm [LM79, CP07] and the Loris–Verhoeven algorithm [LV11] (also known as proximal alternating predictor corrector algorithm). For the three-term sum they include primal-dual fixed point algorithm (PDFP) [CHZ16], primal-dual Davis–Yin algorithm (PDDY) [DY17] and the asymmetric forward-backward adjoint splitting (AFBA) [LP17]. The existing two- and three-term algorithm are often intimately connected. For example, with specific parameter and function choices PD3O is a generalisation of Loris–Verhoeven, Davis–Yin, PDHG and Douglas–Rachford algorithms. We refer to [JV23, CKCH23] for further connection between these algorithms. ADMM can also be extended to solve three-term sum problems, however, as shown in [CHYY16], the resulting algorithm is not necessarily convergent, and thus requires modifications.
Numerical approaches for optimisation problems with more than three terms (sometimes called the multi-block problem) have also been studied in the literature. In [RFP13] the authors proposed a so-called “generalised Forward-Backward splitting” algorithm for the minimisation of a smooth function and several prox-friendly terms. There are several primal-dual splitting methods for solving multi-block problems, e.g. [CP12, Vũ13, CV14]. A widely adopted strategy is block-wise updates, which shares the same spirit as coordinate descent; see also the discussion in Remark 4.
2.3   Further Reading
Here we discuss advanced topics related to presented approaches, including: higher-order schemes, Bregman methods, incremental algorithms and so on.
Second-Order Methods
First-order methods have several drawbacks. For example, they often exhibit slow convergence speed, are not scale invariant, and often require tedious parameter tuning. Higher-order methods such as Newton’s method address some of the issues by using higher-order derivatives to estimate the objective [NW06]. In many imaging applications, computing the Hessian can be prohibitively expensive. Quasi-Newton methods try to overcome this problem by approximating the Hessian using only gradient information. Perhaps the most-known quasi-Newton method is the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [Bro70, Fle70, Gol70, Sha70]. BFGS achieves superlinear convergence under suitable convexity and smoothness conditions, and it has been applied to nonsmooth problems [LO13, YVGS08]. However, its memory requirements can be high, which has led to the development of the limited-memory BFGS (L-BFGS) [LN89], that has been applied to several inverse problems [Vog00, TBE+18].
The superlinear convergence of Newton methods can be extended to nonsmooth optimisation problems for the class of so called semi-smooth functions [QS93]. Semi-smooth Newton methods have been applied for Lasso problems [LST18], and have been extended to PGD and Douglas–Rachford splitting type algorithms [XLWZ18, HTPW22, LST23], primal-dual algorithms [WFO23], to name a few, with applications in TV minimisation [NQYH07] and seismic tomography [BU15].
Bregman Proximal Methods
Most of this review is restricted to the Hilbert space / Euclidean setting. Generalisations are an active field of research, often based on the Bregman distance [Bre67]. Let $\phi:\mathcal{X}\to\mathbb{R}$ be a strictly convex and continuously differentiable function. The Bregman distance associated to $\phi$ is defined as $D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle$. The Bregman distance is in general not symmetric and does not satisfy the triangle inequality. However, it is non-negative and only equals to 00 if and only if $x=y$. The notion of $L$-smoothness can be generalised to “relatively smooth” associated to a Bregman distance [BBT17, Teb18]. Various Bregman first-order methods can be designed in this framework, such as Bregman proximal algorithms [BBT17], Bregman ADMM [WB14], Bregman PGD [MO19, HRX21], and Bregman Primal-Dual splitting [CC22, JV22].
Nonconvex Optimisation
While many applications can be well modelled with a linear forward model, there are also many applications where a linear model is insufficient, e.g. electrical impedance tomography (EIT) [CIN99], phase retrieval for imaging [SEC+15], photoacoustic tomography [WA11], full waveform inversion [Tar84] with applications in geophysics, [VO09], optical imaging generally [AH97] and diffuse optimal tomography specifically [BBM+01], to name just a few. From an optimisation perspective this is challenging as the resulting minimisation problem is generally nonconvex, its regularity properties are heterogeneous and most algorithms are only guaranteed to converge to locally optimal solutions. Some convex optimisation algorithms have been extended to nonconvex optimisation and applied to inverse problems [FG98, CGOT00, Val14, Val21, CZX+21].
Learning-to-Optimise
With the emergence of deep learning, learning-to-optimise (L2O) is becoming a promising direction in optimisation [LM16, CCC+22]. This line of work was in its early days aimed at accelerating the training of machine learning models, focusing mostly on improving the network training, compared to classical hand-crafted optimisers such as SGD and Adam using reinforcement learning [ADG+16]. However, most such methods lack convergence guarantees which motivated research on provably convergent L2O schemes, which aim to learn parts of the algorithm such as the step-sizes for PDHG [BRA+20] and the mirror maps of the mirror descent algorithms [TMTS23, TMTS24a], with applications in both machine learning and inverse imaging problems. In [EFG24], the authors developed a framework for greedy L2O with guaranteed convergence on the training data. Theoretical results of the generalisation guarantees for L2O schemes were recently developed [SFO24], leading a new dimension of research in this area.
3   Stochastic Algorithms for Variational Regularisation
The first-order numerical schemes discussed in Section 2 have been the predominant methods for solving inverse problems over the past decades, due to their efficiency, simplicity, and relatively fast convergence speed. However, due to rapid increases in problem complexity and data sizes over the recent years, the computational costs associated with first-order methods is becoming prohibitively high. This is the main driving force for the development of algorithms with a low per-iteration computational cost, which are often stochastic in nature. In this section we present an overview of existing stochastic optimisation approaches particularly relevant for inverse problems. The layout of the section is similar to that of Section 2. In Section 3.1 we discuss the basic ingredients of randomisation for first-order optimisation algorithms, and in Sections 3.2 and 3.3 we introduce several representative algorithms. We conclude in Section 3.4.
3.1   Designing First-Order Stochastic Algorithms
When designing stochastic optimisation algorithms we aim to reduce the per-iteration computational costs while not degrading the performance too much, thus creating an overall faster and more efficient approach. In order to discuss the design principles of stochastic algorithms, we first rewrite problem (7), revealing a finite-sum structure,
$\min_{x\in\mathcal{X}}\left\{\Phi(x)=\sum_{i=1}^{\ell}f_{i}(A_{i}x)+\sum_{i=1}^{m}g(x)+\sum_{i=1}^{n}h_{i}(x)\right\}.$ (22)
For all $i$ the operators $A_{i}:\mathcal{X}\rightarrow\mathcal{Y}_{i}$ are linear and bounded, $f_{i}:\mathcal{Y}_{i}\rightarrow\mathbb{R}_{\infty}$ and $g_{i}:\mathcal{X}\rightarrow\mathbb{R}_{\infty}$ are proper, closed, convex, and prox-friendly, and $h_{i}:\mathcal{X}\rightarrow\mathbb{R}$ are convex and $L_{i}$-smooth. Moreover, we denote $\mathcal{Y}=\mathcal{Y}_{1}\times\ldots\times\mathcal{Y}_{\ell}$. The analysis of stochastic algorithms for this problem relies on several key quantities, the most used one is the largest Lipschitz constant of the smooth terms, $L_{\max}=\max\{L_{1},\ldots,L_{n}\}$.
Remark 6
The finite-sum structure in (22) naturally arises in inverse problems when considering MAP estimators, see Remark 1. In contrast, in machine learning, the loss function often represents an empirical risk which consists of the average loss, i.e. each sum is divided by the number of summands. While mathematically equivalent, this rescaling has consequences for some formulas that follow.
Example 3.1
Let $K$ be a forward operator, modelled by the discrete X-Ray transform with a parallel-beam geometry and $180$ angles from the $[0,\pi)$ range, $u^{\dagger}$ the object of interest, and $v$ the observed sinogram (encoding the noise as well). The forward operator $K$, object $u^{\dagger}$, and data $v$ are approximately related via the forward model $Ku^{\dagger}=v$. To recover $u^{\dagger}$ from $v$ we consider the variational problem
$\operatornamewithlimits{argmin}_{x\in\mathcal{X}}\Big{\{}\Phi(x)=\frac{1}{2}\|Kx-v\|_{2}^{2}+\alpha\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)\Big{\}},$ (23)
where $\alpha>0$. The objective in (23) can be written in several ways to deal with the costs of evaluating $K$ and $K^{\ast}$. For the following, let $K_{i}$, for $1\leq i\leq 30$ be the CT Radon operator that uses $6$ angles from the range $[\pi i/180,\pi)$ with increments of size $\pi/30$. The data $v$ is split in a corresponding fashion, giving subset data vectors $v_{i}$ that use $1/30$-th of the data in $v$. We can then identify functions in (23) with the template (22) as
  • $g(x)=\alpha\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $f\equiv 0$, and $h_{i}(x)=\frac{1}{2}\|K_{i}x-v_{i}\|_{2}^{2}$, $n=30$;
  • $g(x)=\alpha\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $A_{i}=K_{i}$ with $f_{i}(x)=\frac{1}{2}\|x-v_{i}\|^{2}$, $\ell=30$ and $h\equiv 0$.
As in Section 1, matching the inverse problem to the template is not unique and an algorithm design question. In Section 5 we will revisit this example numerically.
Similar to Section 2.1, we first discuss basic approaches for addressing these challenges and ideas to design stochastic algorithms.
3.1.1   Basic Ingredients
Stochastic Gradient Approximation
The underlying idea of stochastic gradient methods is to approximate the full gradient of the smooth term by a stochastic estimator. The prototypical example is the stochastic gradient descent (SGD) [RM51] which in each step approximates the full gradient $\nabla h(x^{(k)})=\sum_{i=1}^{n}\nabla h_{i}(x^{(k)})$ by a gradient of a single function ~$\tilde{\nabla}h(x^{(k)})\approx n\nabla h_{i_{k}}(x^{(k)})$, using a uniformly at random selected index $i_{k}$. As a result, the per-iteration computational effort is reduced by a factor of $1/n$, and thus one iteration of GD corresponds to executing $n$ iterations of SGD. This is referred to as one epoch or data pass. The drawback of SGD is that the error introduced by the variance of the random process is not reduced along the iterations. A zoo of algorithms has been developed to address this issue over the years, which together with SGD, will be discussed in more detail in subsection 3.2.
Stochastic Proximal Point Algorithms
The stochastic proximal point algorithm is the stochastic counterpart of the PPA for non-smooth problems, by evaluating a single randomly chosen proximal operator only, i.e. $x^{(k+1)}=\operatorname{prox}_{\tau_{k}g_{i_{k}}}(x^{(k)})$ [Bia16]. Similarly to the smooth setting, variance reduced variants can be derived [TASV23].
Randomised Dual Coordinate Descent
As is the case for deterministic algorithms, stochastic algorithms can also be applied to the dual formulation of the problem. Namely, the dual problem to $\min_{x}\sum_{i=1}^{\ell}f_{i}(A_{i}x)$ is computed as $\min_{y}\sum_{i=1}^{\ell}f_{i}^{*}(y_{i})$ subject to $\sum_{i=1}^{\ell}A_{i}y_{i}=0$ [SSZ13]. Notice that the finite-sum structure in the primal problem leads to a dual problem which is separable in $y$ and thus naturally amendable to dual coordinate descent strategies [RT14]. This idea has been exploited in primal-dual algorithms for (22) like QUARTZ [QRZ15] and SPDHG [CERS18]. It is worth noting that stochastic gradient descent and stochastic coordinate descent are deeply connected via duality, and we refer to [SSZ13, GR15] for discussions.
Table 2 summarises the basic ingredients for the design of first-order stochastic optimisation algorithms. Next we provide detailed explanations of these methods and discuss specific state-of-the-art algorithms that use these concepts.
Target Tool References
$\sum_{i=1}^{n}h_{i}(x)$ Stochastic gradient approximation [RM51, BCN18]
$\sum_{i=1}^{n}h_{i}(x)$ Better stochastic gradients with history [DBLJ14, JZ13]
$\sum_{i=1}^{m}g_{i}(x)$ Stochastic proximal operators [Bia16, TASV23]
$\sum_{i=1}^{\ell}f_{i}(A_{i}x)$ Coordinate descent for dual problem [SSZ13, CERS18]
Table 2: Elements in designing stochastic first-order algorithms.
3.1.2   Designing Stochastic Algorithms
Once the computational expensive components of a deterministic optimisation algorithm have been identified, they can be replaced with a stochastic alternative. However, there are still several delicate tasks, for example, decomposing (7) into the finite-sum form (22) is not unique, choosing $\ell,m,n$ and partitioning $A$ such that the performance of the algorithms are optimised is not a straightforward task in general.
In the rest of the section we introduce several representative stochastic optimisation algorithms and discuss related topics.
3.2   Stochastic Gradient Methods
In this section we review in detail algorithms based on stochastic approximation of the gradient, which dates back to 1950s. In their seminal work [RM51], Robbins and Monro realised that data are often highly correlated and exhibit a large degree of redundancy, thus using all available data in each computation is unnecessary and inefficient. Another insight is that GD is quite robust in the sense that convergence towards the minimum can be ensured even when there are small errors in the descent direction, as long as the errors decrease in the long run. This has significant practical benefits, allowing to efficiently solve large scale optimisation problems.
Since the 2010s, stochastic gradient schemes have seen a resurgence of interest, due to the machine learning renaissance, and the library of existing methods has been significantly enriched. The new algorithmic modifications include adaptive steps-sizes, variance reduction schemes, and many more.
3.2.1   Vanilla Stochastic Gradient Descent
When the composite terms vanish and we consider a single prox-friendly function, i.e. $\ell=0$, $m=1$, template (22) reduces to
$\min_{x\in\mathcal{X}}\left\{\Phi(x)=g(x)+\sum_{i=1}^{n}h_{i}(x)\right\}.$ (24)
Stochastic gradient descent (SGD) [RM51] is a classical stochastic optimisation method for efficiently solving this type of a problem, see Algorithm 7. In each iteration $k$ it performs a step in the (negative) direction of $\nabla h_{i_{k}}(x)$ for a randomly sampled index $i_{k}\in\{1,\ldots,n\}$ using a pre-defined step-size schedule $\tau_{k}>0$ (a.k.a. learning rate).
Input: $x^{(0)}\in\mathcal{X}$, $\tau_{k}>0$.
1 for $k=0,1,\ldots$ do
2       Sample $i_{k}\in\{1,\ldots,n\}$ uniformly at random
3       ~$\tilde{\nabla}h(x^{(k)})=n\nabla h_{i_{k}}(x^{(k)})$
4       $x^{(k+1)}=\mathrm{prox}_{\tau_{k}g}\big{(}x^{(k)}-\tau_{k}\tilde{\nabla}h(x^{(k)})\big{)}$
5 end for
Algorithm 7 Stochastic Gradient Descent
Instead of using a single function $h_{i}$ in each iteration, in machine learning it is common to use several randomly selected functions, resulting in the so-called mini-batch SGD. In contrast, for inverse imaging problems it is more common to pre-batch the data by partitioning the forward operator $K$, and define functions $h_{i}$ from the partition. In this context, batches are usually referred to as subsets, and thus the number of batches and batch size become number of subsets and subset size. We will discussed subset selection in more detail in Section 4. SGD has been used in image reconstruction, e.g. for PET [TAK+23], CT [GB18], optical tomography [CLL18] and cryo-EM [SS20, PRFB17].
The descent direction in SGD is an unbiased estimator of the full gradient, that is, $\mathbb{E}[n\nabla h_{i}(x)]=\nabla h(x)$. A common assumption needed to show the convergence is the bounded variance property $\mathbb{E}[\|n\nabla h_{i}(x)-\nabla h(x)\|^{2}]\leq C,$ for some $C>0$. The variability between successive iterates is a critical issue for SGD. SGD does not converge to the minimum for constant step-sizes. Instead, vanishing step-sizes are needed to control the variability [RM51]
$\sum_{k=1}^{\infty}\tau_{k}=\infty,\quad\sum_{k=1}^{\infty}\tau_{k}^{2}<\infty.$ (25)
For example $\tau_{k}=k^{-\delta},\delta\in(1/2,1]$ satisfies the above conditions, though other choices are also popular, e.g. [BCN18]. However, decaying step-size regimes eventually lead to a very slow speed of convergence. To circumvent the slowdown in convergence, there are in general two types of modifications: variance reduction schemes and adaptive step-sizes.
Remark 7
In Algorithm 7 indices $i_{k}$ are sampled at random, which is typically done according to a uniform distribution in the discrete set $\{1,\ldots,n\}$, without replacement. Alternatively, a probability $p_{i}$ is assigned to each $i\in\{1,\ldots,n\}$. Choice of the sampling probability distribution can reduce the variance in SGD. However, for certain problems it can be shown that the optimal probability distribution is adaptive and requires solving a minimisation problem (in each iteration) that can be harder than the original problem. For $L$-smooth functions setting $p_{i}=L_{i}/\sum_{j=1}^{n}L_{j}$ is a good proxy [ZZ15].
Remark 8
As previously discussed, SGD is more commonly applied in machine learning where it is used for empirical risk minimisation, giving an averaged objective, which is not the case in inverse problem, see Remark 1. Consequently, due to the lack of a multiplicative constant in front of the summed terms, the stochastic direction in Algorithm 7 is scaled with respect to the number of functions.
This has further desirable consequences. First, scaling with $n$ ensures that the update direction is of the same order of magnitude as the full gradient, which is also true when SGD is applied in empirical risk minimisation. Second, from a probabilistic perspective scaling with $n$ is needed to ensure that the descent direction is an unbiased estimator of the full gradient.
3.2.2   Variance Reduction Schemes
Stochastic variance reduction algorithms aim to reduce the variance present in SGD through historical gradient information. In the traditional statistics literature such quantities are sometimes referred to as control variates. These approaches combine the low-per iteration complexity of SGD with constant step-sizes (and faster convergence rates) of PGD, at a cost of increased storage requirements. The study of variance reduction algorithms began with Stochastic Average Gradient (SAG) [RSB12], Stochastic Averaged Gradient Amélioré [^5] The term amélioré is used in [GSBR20]. It is French and translates to English as “improved”. (SAGA) [DBLJ14], and Stochastic Variance-Reduced Gradient (SVRG) [JZ13], but has since grown significantly. We refer to [DLS22, GSBR20] for an overview. Variance reduction algorithms have been used for many image reconstruction problems, such as PET [TAK+23, KTA+21], CT [KW17], optical tomography [MAP20], seismic tomography [HCHY23] and blind deblurring [DTL+21]. Below we describe SAGA and SVRG, two representative variance-reduced variants of SGD.
SAGA and SAG
In order the reduce the variance, SAG and SAGA imitate PGD by storing a table of the latest computed gradients for each sub-function $h_{i}$. The entire table is used in each iteration as the approximation of the full gradient, and thus information from all the gradients $\nabla h_{i}$, which lowers the variance. However, the table is updated only at the position of the most recently used (i.e. randomly sampled) function index, which reduces the computational effort.
Input: $x^{(0)}\in\mathcal{X}$, $\tau>0$
1 for $k=0,1,\ldots$ do
2       Sample $i_{k}\in\{1,\ldots,n\}$ uniformly at random
3       ~$\tilde{\nabla}h(x^{(k)})=n\big{(}\nabla h_{i_{k}}(x^{(k)})-\gamma_{i_{k}}^{(k)}\big{)}+\sum_{i=1}^{n}\gamma_{i}^{(k)}$
4       $x^{(k+1)}=\operatorname{prox}_{\tau g}\big{(}x^{(k)}-\tau\tilde{\nabla}h(x^{(k)})\big{)}$
5       $\gamma_{j}^{(k+1)}=\begin{cases}\nabla h_{j}(x^{(k)})&\textrm{if~{}~{}}j=i_{k},\\ \gamma_{j}^{(k)}&\textrm{otherwise}\end{cases}$
6 end for
Algorithm 8 Stochastic Average Gradient Amélioré (SAGA)
When $h_{i}$ are convex, the average iterate of SAGA converges to a minimiser of the objective at a sublinear rate $\mathcal{O}(1/k)$ provided $\tau=1/(3nL_{\max})$ [Def15]. Note that the convergence of SAGA is given in terms of the average iterate, and not the final iterate, as common for deterministic algorithms in Section 2. This is a fairly common issue for stochastic optimisation algorithms, and arises due to proof techniques. However, in practice it is common to use the final iterate instead of the average one, regardless of the missing theoretical guarantees.
SAG follows nearly identical iterations, except that there is no $n$ factor in front of $\nabla h_{i_{k}}(x^{(k)})-\gamma_{i_{k}}^{(k)}$, which lowers the variance and makes SAG biased.
The original convergence analysis of SAG required a computer aided approach to complete the proof, though the convergence was recently proven without the need for computer aid [DLS22].
Storing the table of historical gradients can become infeasible when the number of subsets is large, or when data is of a high dimensionality. However, when applied to linear models (in the sense of [GSBR20]), the memory requirement can be significantly improved in certain applications, see Section 4 for details.
Remark 9
Momentum based acceleration methods can be seen as a form of variance reduction, as they use a weighted average of previous gradients which gives more weight to recent gradients. In contrast, SAG(A) uses a plain average of the recent gradients.
SVRG
Instead of storing a table of gradients, SVRG stores a reference iterate, often called an anchor point, or a snapshot, as well as the full gradient evaluated at the reference iterate. Both the reference iterate and the full gradient are updated at pre-defined intervals during the iterations. To this end, SVRG requires two loops: in the outer loop we update the reference iterate and compute the full gradient, and in the inner loop we perform stochastic updates using a combination of the current gradient and the full gradient from the outer loop. The length $t\in\mathbb{N}$ of the inner loop is by convention set to $2n$ for convex and $5n$ for nonconvex problems.
Input: $x^{(0)}\in\mathcal{X}$, $\tau>0$, inner loop length $t\in\mathbb{N}$
1 Set $\overline{x}^{(0)}=x^{(0)}$
2 for $j=0,1,\ldots$ do
3       Compute $\overline{\gamma}^{(j)}=\sum\nolimits_{i=1}^{\ell}\nabla h_{i}(\overline{x}^{(j)})$
4       for $k=0,\ldots,t-1$ do
5             Sample $i_{k}\in\{1,\ldots,n\}$ uniformly at random
6             ~$\tilde{\nabla}h(x^{(k)})=n\big{(}\nabla h_{i_{k}}(x^{(k)})-\nabla h_{i_{k}}(\overline{x}^{(j)})\big{)}+\overline{\gamma}^{(j)}$
7             $x^{(k+1)}=\operatorname{prox}_{\tau g}\big{(}x^{(k)}-\tau\tilde{\nabla}h(x^{(k)})\big{)}$
8       end for
9      Set $\overline{x}^{(j+1)}=x^{(t)}$
10 end for
Algorithm 9 Stochastic Variance Reduced Gradient (SVRG)
SVRG was initially designed only for smooth objectives [JZ13]. The proximal version for composite objectives of the type (24) was developed and analysed in [XZ14]. They show for strongly convex problems that the average iterate of SVRG converges in the objective value at a linear rate for $\tau<1/(4nL_{\max})$ and sufficiently large $t$ [XZ14].
The advantage of SVRG over SAG(A) is that it requires storing only two pieces of data, i.e. $\overline{x}^{(j)}$ and $\overline{\gamma}^{(j)}$, hence its memory requirements are independent of $n$. However, its disadvantages are the need for two loops, which introduces an additional tuning parameter, and the need to evaluate the full gradient every time the reference iterate is updated (which is not needed by SAGA).
Adaptive Subset Sizes
The number $n$ of smooth terms, or equivalently the batch size, has a direct impact on the variance of SGD and can be derived for general cases of unbiased stochastic estimators with bounded variance [BCN18]. Specifically, the variance of the stochastic estimator increases as the number of subsets increases, at a roughly linear rate. Therefore, variance can be reduced not by reducing the step-size but rather by adaptively or dynamically selecting the batch or subset size. This topic has been predominantly studied from the deep learning perspective, e.g. [SKL18, DNG17, BRH17]. However, it is important to note that increasing $n$ also increases the per-iteration computational cost and can thus run into same scalability issues that PGD encounters. Thus, timely increases of the batch size are critical in practice, and often specific criteria are designed to address this issue, e.g. [BCNW12]. While different variations of this methodology have been used in certain inverse problems, such as diffuse optical tomography [MAP20] and PET [TAH+19], the topic has not yet seen widespread adoption.
3.2.3   Further Stochastic Gradient-Based Methods
SGD and variance reduction algorithms have in recent years seen many modifications and adaptations to specific scenarios of interest. For example, Loopless-SVRG algorithm [KHR20] is a modification of SVRG that uses only one loop and instead of updating the reference iterate and full gradient at pre-defined intervals, in each iteration it takes a random sample from $[0,1]$. If the sampled value is below a predefined probability, reference and full gradient are updated, and otherwise a standard update is perform. SARAH [NLST17] is a biased variant of SVRG that updates the full gradient not only in the outer loop but also in the inner loop. Similar biased variance-reduce gradient scheme includes SPIDER [FLLZ18, WJZ+19]. See also [DLS22] for a gradient estimator named SARGE which combines the ideas of both SAGA and SVRG.
Algorithms for the three-term sum from Section 2.2.2 also admit stochastic variants. For example, stochastic variants of PD3O or Condat–Vũ algorithms can be combined with SGD or SVRG in a mix-and-match fashion by replacing the gradient $\nabla h(x)$, with a corresponding stochastic estimator, see [SCMR22]. Using stochasticity in the dual variable is discussed in Section 3.3.
Principles used for SGD, SAGA, and SVRG have also been applied to maximum likelihood expectation maximisation, and MAP, see [CM09, CZTZ18, KWML19], and have been applied to tomographic image reconstruction [KTA+21].
3.2.4   Accelerated Stochastic Variance-Reduced Gradient Methods
As variance-reduced gradient methods can use constant step-size and have the same convergence rate as their deterministic counterparts, a natural question to ask is whether NAG/FISTA can be extended to them. This is a rather challenging task as NAG/FISTA are very sensitive to errors [AD15].
The first attempt along this direction is [Nit14], in which FISTA is combined with SVRG gradient estimator, however no theoretical guarantee when the problem is not strongly convex.
Later on in [LMH15], based on an accelerated proximal point algorithm, a so-called Catalyst acceleration is proposed to speeding up the solving of (24). Catalyst contains two loops, the outer loop is accelerated proximal point algorithm, while the inner loop is solved with a stochastic gradient scheme to certain accuracy. In [AZ17], an acceleration scheme named Katyusha is proposed for accelerating SAGA and SVRG. By incorporating a “negative momentum” term, the author manages to prove an $\mathcal{O}(1/k^{2})$ convergence rate without strong convexity. Later on in [DES22], a more elegant acceleration scheme is proposed, see Algorithm 10 below, and the authors show that the negative momentum is not needed to achieve the $\mathcal{O}(1/k^{2})$ convergence.
Input: $x^{(0)},y^{(0)},z^{(0)}\in\mathcal{X}$, $\tau>0$, $\eta_{k}>0$
1 for $k=0,1,\ldots$ do
2       $x^{(k+1)}=\eta_{k}z^{(k)}+(1-\eta_{k})y^{(k)}$
3       Compute a variance-reduced stochastic gradient estimate ~$\tilde{\nabla}h(x^{(k+1)})$
4       $z^{(k+1)}=\operatorname{prox}_{\tau g}\big{(}z^{(k)}-\tau\tilde{\nabla}h(x^{(k+1)})\big{)}$
5       $y^{(k+1)}=\eta_{k}z^{(k+1)}+(1-\eta_{k})y^{(k)}$
6 end for
Algorithm 10 Universal Scheme for Accelerating Variance-Reduced Methods
In recent years, vector extrapolation schemes are also adopted in the stochastic optimisation, for example the nonlinear acceleration [SBd17] schemes (e.g. the Anderson Acceleration and its variants) are applied to accelerate the stochastic gradient methods. This type of schemes are very closely related to quasi-Newton schemes such as L-BFGS, as showed in [Sci19].
3.2.5   Incremental Gradient Methods
Before wave of stochastic optimisation algorithms for machine learning in the 2010s, another related family of optimisation algorithms received increased interest: incremental gradient methods. Incremental gradient methods are also designed for solving objective functions with the finite-sum structure (24) as for SGD (see Algorithm 7). The core idea is that in each step of the iteration, the algorithm chooses one function from the finite-sum via a predetermined rule (cyclic or fixed permutation), and apply GD to the chosen function [Ber11]. Similar to sub-gradient descent and SGD, to ensure convergence, a carefully selected step-size is required: $\sum_{k}\tau_{k}=\infty$, $\sum_{k}\tau_{k}^{2}<\infty$. This idea is used in the modified block sequential regularised expectation-maximisation (BSREM) and the relaxed OS separable paraboloidal surrogates (OS-SPS) [PY01, AF03], the former is used currently in clinical PET imaging [SSD+17].
To circumvent the need for a diminishing step-size, an incremental averaged gradient (IAG) scheme was developed in [BHG07] which can afford constant step-size. The basic idea of the method is averaging the history gradients, which is also the key behind stochastic averaged gradient methods SAG and SAGA (see e.g. Algorithm 8[SLRB17, DBLJ14].
This methodology can be extended to non-smooth optimisation (e.g. for PGD). For instance when the non-smooth function is decomposable into a finite-sum, one can choose a single term in the sum and compute its proximal operator. We refer to the dedicated survey [Ber11] for more details.
Lastly, compared to the stochastic counterpart, incremental gradient methods have weaker theoretical guarantees, such as more restrictive step-size criteria [Ber11, BHG07, SLRB17]. The numerical investigation of [TAK+23] in the context of PET imaging showed that BSREM was inferior to stochastic variance reduced algorithms in every considered metric. See Section 5 for a comparison of incremental gradient methods to stochastic methods. We add that incremental gradient algorithms that access gradient indices in a cyclic manner can show exponentially bad behaviour for certain pathological use-cases [HM93, RR12, Def15].
3.3   Stochastic (Dual) Coordinate Descent
In the context of inverse problems, stochastic gradient methods often reduce the complexity in measurement space. For example, when the smooth term $h$ corresponds to data fidelity, e.g. $\frac{1}{2}\|Kx-v\|^{2}$. An alternative is to reduce the complexity in the space of the optimisation variable. Recall the CD from Section 2.1.1, which performs the iterations
$x^{(k+1)}=x^{(k)}-\tau_{k}\nabla_{i}h(x^{(k)}),$
where $\nabla_{i}h(x)=\frac{\partial h(x)}{\partial x_{i}}e_{i}$ and $e_{i}$ is the $j$-th canonical basis vector. As in the deterministic case, instead of using a single coordinate we can use a subset of (randomly selected) coordinates. Note that the direction $\nabla_{i}h(x)$ is an unbiased estimation of $\nabla h(x)$. Moreover, it is also variance-reduced. If $x^{(k)}\rightarrow x^{\ast}$, then due to the optimality condition of $x^{\ast}$, we get $\nabla h(x^{(k)})\to 0$ and $\nabla_{i}h(x^{(k)})\to 0$, thus $\|\nabla_{i}h(x^{(k)})-\nabla h(x^{\ast})\|^{2}\rightarrow 0$.
To provide a reduction in computational costs it usually requires coordinate-friendly structures (such as separable functions), that is, coordinate updates should be significantly cheaper than gradient computations. Because the optimisation problems associated with inverse problems usually do not exhibit this structure, coordinate descent methods are largely not used in this context. An exception is for instance [GB18] which uses coordinate descent in large-scale CT image reconstruction to overcome memory bottlenecks on GPUs. It turns out that coordinate descent for inverse problems is more useful when applied on dual variables as we will see below.
Whereas stochasticity has been observed to improve the performance for gradient estimators, as compared to using a deterministic (e.g. cyclic) sampling of the gradients, the situation for coordinate descent is more nuanced. On one hand, there are classes of non-convex problems for which cyclic CD fails to converge to a stationary point [Pow71], but stochastic CS (SCD) is expected to converge within only a few steps, and in other cases SCD can be significantly faster than cyclic CD [Wri15]. On the other hand, one can design classes of forward operators such that cyclic CD for the corresponding quadratic problem converges several times faster than SCD [GOPV17]. For a further comparison between cyclic, stochastic, and adaptive choices for block CD we refer to [NLS22].
3.3.1   Stochastic Primal-Dual Splitting
For simplicity we will restrict ourselves to the case when $m=1$ and $n=0$, i.e. our template (22) takes the form
$\min_{x\in\mathcal{X}}\left\{\Phi(x)=\sum_{i=1}^{\ell}f(A_{i}x)+g(x)\right\},$ (26)
with associated saddle-point problem
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\left\{\sum_{i=1}^{\ell}\left(\langle A_{i}x,y_{i}\rangle-f_{i}^{*}(y_{i})\right)+g(x)\right\}.$ (27)
When applying PDHG (Algorithm 3) to this problem every iteration requires evaluating matrix-vector products for all $A_{i}^{*}$ to update the primal variable and $A_{i}^{\ast}$ to update the dual variable which can be computationally costly. Thus, in [CERS18] it was proposed to perform a dual coordinate ascent instead. The resulting algorithm is called Stochastic Primal-Dual Hybrid Gradient (SPDHG) and provided in in Algorithm 11.
Input: $x^{(0)}\in\mathcal{X},y^{(0)}=\overline{y}^{(0)}\in\mathcal{Y}$, $\tau,\sigma>0$
1 for $k=0,1,\ldots$ do
2       $x^{(k+1)}=\operatorname{prox}_{\tau g}(x^{(k)}-\tau A^{*}\overline{y}^{(k)})$
3       Sample $i_{k}\in\{1,\ldots,\ell\}$ uniformly at random
4       $y^{(k+1)}_{i}=\begin{cases}\operatorname{prox}_{\sigma f_{i}^{*}}(y^{(k)}_{i}+\sigma A_{i}x^{(k)})&\text{if $i=i_{k}$}\\ y^{(k)}_{i}&\text{else}\end{cases}$
5       $\overline{y}^{(k+1)}_{i}=\begin{cases}y^{(k+1)}_{i}+\ell(y^{(k+1)}_{i}-y^{(k)}_{i})&\text{if $i=i_{k}$}\\ y^{(k)}_{i}&\text{else}\end{cases}$
6 end for
Algorithm 11 Stochastic Primal-Dual Hybrid Gradient (SPDHG)
SPDHG has been shown to converge in Bregman distance if $\tau\sigma<1/(\ell\max_{i}\|A_{i}\|^{2})$. This convergence result can be extended and the algorithm be generalised to other samplings and step-sizes, see [CERS18, AFC22, GDE21] for more details.
Note that due to the linearity of $A$, we can rewrite the primal update as
$A^{*}\overline{y}^{(k+1)}=(1+\ell)A^{*}_{i_{k}}(y^{(k+1)}_{i}-y^{(k)}_{i})+\sum_{i=1}^{\ell}A^{*}_{i}y^{(k)}_{i}.$
Note that this bears a close resemblance to a SAGA estimator of the gradient of the bilinear term in (27) with respect to $x$. However, due to the factor $1+\ell$ instead of $\ell$, this is a biased gradient estimator. Its effectiveness on large-scale problems has been shown in various applications such as PET with sinogram data [EMS19], PET with listmode data [SH22], CT [CDE+24, RDH21, PAD+21], multi-spectral CT [PAD+21], parallel MRI [GDE21, GDE22] and magnetic particle imaging [ZB21].
As for any primal-dual algorithm, it is difficult to tune the ratio of the primal and dual step-sizes. This is addressed empirically in [ZB21] and with theoretical foundations provided in [CDE+24]. Note that the problem can also be approached similarly by other primal-dual algorithms, e.g. [SSZ13, FR15, FB19].
3.4   Further Reading
Stochastic Second-Order Methods
Standard second-order methods have seen limited application in large scale problems, due to the costs and intractability in computing the Hessian of the objective. As discussed in Section 2.3, deterministic quasi-Newton methods use information about the curvature without computing the Hessians. Instead, approximation techniques are used to approximate the Hessian or the corresponding matrix-vector products. A direct extension of quasi-Newton methods to the stochastic setting can be challenging, since noisy estimates of the curvature (coupled with stochastic estimators of the gradient) can destabilise the iterations. A number of methods have been proposed to solve this issue, and leverage the strengths of quasi-Newton and stochastic gradient methods.
A pioneering work in this field is the stochastic quasi-Newton method [SYG07], which introduced stochastic modifications of BFGS and L-BFGS, showing a sub-linear convergence rate. This limitation was addressed by [MNJ16], achieving linear convergence rates under mild assumptions, through a combination of the stochastic variant of L-BFGS [BHNS16] with variance reduction.
Sub-sampled quasi-Newton methods [RKDA14, EM15, XYR+16, RKM19] randomly sub-sample the Hessian to further reduce the costs, and use matrix concentration inequalities to ensure that information about the curvature is well preserved. Despite these advancements, stochastic quasi-Newton methods are still not widely adopted neither in the inverse imaging community, with notable counterexamples [MXC+19, HPAT22, PA21] and they continue to be an active area of research.
Forward Operator Sketching
The underlying idea of sketching methods is to replace the original inverse problem with an inverse problem that is much smaller and thus more manageable. This is often studied for the least-squares problem $\min_{x\in\mathbb{R}^{d}}\|Kx-v\|^{2}$. The idea is to find a sketching matrix $S\in\mathbb{R}^{m\times s}$, where $m\ll s$ such that the solutions of
$\min_{x\in\mathbb{R}^{d}}\|S(Kx-v)\|^{2}$
are, in a certain sense, similar to the solutions of the original problem, but are drastically cheaper to compute.
Common sketching techniques include random projections, sub-sampling, and leverage score sampling [Woo14, DMM06, Mah11]. This type of sketching comes at a cost since there is an irreversible loss of information [PW16].
Iterative Hessian Sketching
An alternative way to perform sketching for inverse problems is via iterative Hessian sketching (IHS). As Newton’s method, it construct quadratic approximations to the original problem via Hessians which are then replaced by sketched versions of it. In more detail, consider minimising a smooth convex objective $h(x)$, at each iteration, a sketching matrix $S\in\mathbb{R}^{m\times d}$ is used to construct a sketched Hessian matrix $H_{S}=(H^{1/2})^{*}S^{*}SH^{1/2}$, where $H$ is the original Hessian matrix at the current iterate $x^{(k)}$ and $H^{1/2}$ is a square-root of it. The sketched Hessian can then be used as
$x^{(k+1)}=x^{(k)}-\tau_{k}H_{S}^{-1}\nabla h(x^{(k)}),$
where $\tau_{k}>0$ is the step-size. For certain convex optimisation problem, IHS is faster than the standard scheme  [PW16, PW17]. It can also be combined with (accelerated) PGD [TGD17] and is applied to accelerate the multi-coil MRI reconstruction [OOI+24].
Particularly common are Gaussian sketches, which have the strongest theoretical guarantees (with respect to compression quality) but are computationally inefficient. They can be accelerated through the Johnson–Lindenstrauss transform (FJLT) [AC09, AL08], and the sparse Johnson–Lindenstrauss transform (SJLT) [CW13], to name a few techniques. In particular, SJLT trades off computational efficiency for theoretical guarantees, but it is often used in practice.
Although they have some clear advantages, sketching-based methods are rarely applied to inverse imaging problems. This can primarily be attributed to the fact that sketching may destroy any beneficial structure present in the forward operator. For example, in CT the Radon transform is a sparse matrix for which we can only use subsampling as the sketching, leading to suboptimal performance. The MRI operator includes FFTs which again prevent naive application of sketching. Recently in [OOI+24], a special sketching operator which can preserve the fast computational structure of FFT is developed and applied to multi-coil MRI reconstruction task.
Stochastic Mirror/Bregman Descent
Stochastic gradient descent has in recent years been extended to the problems that do not rely on the Euclidean structure. Stochastic mirror descent is one powerful such generalisation which in each step can be written in the proximal point form as
$x^{(k+1)}=\operatornamewithlimits{argmin}_{x\in\mathcal{U}}\left\{\tau_{k}\tilde{\gamma}^{k}+D_{\phi}(x,x^{(k)})\right\},$
where $\phi$ is the mirror map and $\tilde{\gamma}^{k}$ is an estimator of the gradient of the objective at $x^{(k)}$. SMD is often investigated through the lens of stochastic schemes for nonsmooth Lipschitz continuous convex functions [CBCG04], though its convergence can be shown for nonsmooth convex objectives [GL12], smooth objectives [Bub15] and other conditions on the objectives, and it has been applied to ill-posed inverse problems [JLZ23]. Adaptive step-size schemes have also been studied in recent years [DLLM23].
SGD has also been designed to when both $\mathcal{V}$ and $\mathcal{U}$ are Banach spaces by applying a proximal point like algorithm [JK23], with applications in CT. In this case selecting $\mathcal{V}$ and $\mathcal{U}$ as certain $\mathcal{L}^{p}$, or Sobolev, spaces can act as an implicit regulariser of the solution and allows adapting to different noise assumptions. However, choosing $\mathcal{V}$ and $\mathcal{U}$ to optimise the reconstruction is a challenging problem. This can to a degree be alleviated by using modular (or variable exponent) Banach spaces [LKEC23], which can adapt to cases of spatially variable noise types.
Adaptive Step-size Methods for Training Neural Networks
In the deep learning era, a significant advance of stochastic gradient descent in the context of training (deep) neural networks is the adaptive step-size stochastic gradient schemes, such as AdaGrad [DHS11], AdaDelta [Zei12], RMSProp [TH12], ADAM [KB15], and so on. We refer to [Rud16] for a brief overview of these methods.
The main motivation behind this type of methods is an observation, that during the training procedure, certain variables are updated much less compared to the others. More precisely, let $d^{(k)}=\tilde{\nabla}h(x^{(k)})$ be the stochastic gradient in each step, then up to $T$ step, if for some index $i$ such that $\sum\nolimits_{k=1}^{T}|d_{i}^{(k)}|$ is much smaller that the accumulation of the rest entries, this results in a much slower training behaviour of $x_{i}$, hence requires longer time for training. To overcome this problem, a basic idea is to set different step-sizes for each entries of $x$. To this end, let $D^{(T)}$ be the accumulation of the stochastic gradient $d^{(k)}$, for example for each $i=1,...,n$, let $D^{(T)}_{i}=\epsilon+\sum_{k=1}^{T}|d_{i}^{(k)}|$ where $\epsilon>0$ is for avoiding the sum is 00. Let $H_{T}=\mathrm{diag}(1/D^{(T)})$ be a diagonal matrix where the diagonal elements is the reverse of $D^{(T)}_{i}$, then an adaptive step-size stochastic gradient descent can take for following step
$x^{(k+1)}=x^{(k)}-\tau H_{k}d^{(k)}.$ (28)
Apparently, if for some $i$ the value $D^{(T)}_{i}$ is very small, then at current step $x^{(k+1)}_{i}$ will have a much larger step-size, hence receives larger update.
Compared to the standard SGD scheme, in general adaptive step-size schemes have two advantages: 1) make tuning of step-size automatic such that each entries converge at relatively same speed; 2) tuning-free step-size choices, for instance in (28) we only need to choose $\tau>0$. Over the past decade, adaptive step-size SGD schemes are more popular in training neural networks, while deeper theoretical understandings are obtained for the standard SGD.
Remark 10
One thing we would like to highlight regarding adaptive step-size stochastic gradient schemes, is that this branch of algorithms is the most dominant choices in machine learning, especially for ADAM algorithm. In the field of inverse (imaging) problems, though the adoption of stochastic algorithms is quite trendy nowadays, adaptive step-size schemes seem received rather limited attention. In comparison, variance reduction schemes are more popular.
Randomised Prox
Algorithms in Section 3 reduce the iteration complexity by using a low-cost estimator of the gradient. Similarly, algorithms in Section 3.3 use a lower-cost update in the dual space. This speeds up the computations in inverse problems since it reduces the computational costs inherent in applying the operator $K$ or $A$. However, the second source of computational costs is in the evaluation of the proximal operator, particularly for algorithms that compute $\operatorname{prox}_{\tau g}$ or $\operatorname{prox}_{\tau f}$ in every iteration. Since stochastic algorithms require more iterations to conduct the same number of data passes (and utilise all of the measurement data), this can incur further computational costs if $\operatorname{prox}_{\tau g}$ is expensive to evaluate.
To address this, new approaches have been developed in recent years that take the stochastic approach to the proximal domain. In [CR23] RandProx algorithm is introduced which uses a stochastic estimator for the proximal step, this refers to the prox of $f^{\ast}$) in the primal-dual Davis–Yin algorithm. The stochastic estimator is in its simplest form defined by a probability parameter $p$, i.e. depending on a coin flip the prox is either applied or not applied. ProxSkip [MMSR22] applies this methodology directly to proximal gradient-style algorithms. However, these methods are predominantly applied to federated or decentralised learning, and their use in inverse imaging problems has not yet been thoroughly investigated.
Stochastic Iterative Regularisation
Regularising property of stochastic optimisers has been a well observed and studied property in the machine learning communities. For linear inverse problems and a least squares type objectives, it has been shown that SGD [JL18, JZZ21] and SVRG [JZZ22] for well chosen a priori stopping criteria have a regularising property. Similar results have been shown for mirror and Banach space variants of SGD [JLZ23, JK23], for randomised Kaczmarz methods [NT14, TL23], and other stochastic optimisation schemes.
Nonlinear Inverse Problems
For some suitably-structured nonlinear inverse problems, the stochastic gradient methods have been successfully applied to. For example the randomised Kaczmarz method and its SVRG-inspired variance reduced variants have been successfully applied to phase-retrieval problems [XLTW22] and EIT [vdDA12, ZQTL24]. However, in general the application of stochastic gradient methods can be more challenging than the inverse problems with a linear forward operators, due to the additional level of difficulty in the optimisation objective. One obvious potential issue of applying stochastic gradient methods on solving nonlinear inverse problem is the difficulty in performing back-tracking line search for step-sizes.
4   Challenges for Stochastic Optimisation for Inverse Problems
In the previous two sections, we have provided an introduction and overview of both deterministic and stochastic optimisation algorithms. In this section we present a thorough discussion on various theoretical and practical aspects of stochastic optimisation methods.
4.1   When and Why Are Stochastic Gradient Methods Fast?
As we mentioned in the beginning of Section 3, the aim of developing stochastic optimisation algorithms is to deal with large-scale optimisation problems. One may naturally wonder are stochastic optimisation methods guaranteed to be faster than their deterministic counterpart? Below are some key points.
Stochastic Acceleration Factor
Whether an inverse problem is suitable or not for the application of stochastic gradient method largely depends on the structure of the optimisation problem. We start by introducing studies on fundamental linear inverse problems which can assist us to have understand better when and why stochastic gradient methods can be efficient/inefficient in different scenarios of inverse problems.
In the works of [TED19, TEGD20], a metric for estimating whether the inverse problem is suitable for applying stochastic gradient methods is proposed. This metric is named stochastic acceleration factor $\Upsilon:=L/L_{\max}$ which is proportional to the ratio of two Lipschitz constants. For the ideal cases, the ratio between the gradient Lipschitz constant is close to $n$. Many machine learning problems with common datasets fall into this category, as well as some important inverse imaging problems such as the CT imaging. However, there are many negative examples with $L_{\max}\approx L$, making $\Upsilon\approx 1$ where the stochastic gradient methods are not recommended. These include compressed sensing inverse problems, image denoising or image inpainting. For image deblurring it strongly depends on the (size of) the kernel, see also the numerical example in Section 5.
In the context of linear inverse problems where the data-fit is chosen to be the least-squares, a specialised analysis reveals that the success of stochastic acceleration is dependant on the structure of the forward operator. The theoretical analysis provided by [TEGD20] showed that an inverse problem is suitable for stochastic gradient methods if and only if the Hessian of the inverse problem has a fast-decaying spectrum. It is often neglected in the optimisation community, that stochastic gradient algorithms are in fact exploiting the intrinsic low-dimensional structure of the data matrices, and if such structure is weak in certain problems, then there will be no benefit using them.
When are Variance-Reduced Methods Beneficial?
The importance of variance-reduced methods depends highly on the amount of noise in the measurement. It has been shown in [VML+19] that SGD converges linearly in the so-called interpolation regime, which in the context of inverse problems essentially means that there is no measurement noise. SGD can be even faster than variance-reduced methods as it can use a constant step-size of $1/(nL)$ in this case [GLQ+19]. In contrast variance-reduced methods typically require smaller constant step-sizes, for instance the largest known provably-convergent step-size is $1/(3nL)$ for SAGA [DBLJ14], while $1/(4nL)$ for SVRG [XZ14]. For the case where the measurement noise is nonnegligible, then the vanilla SGD must reduce its step-size for convergence and become slow for high accuracy solutions. In such cases, variance-reduced stochastic gradient methods are recommended due to their fast convergence in this regime which is prevalent for inverse problems. See also the numerical experiments in Section 5.
Efficient Partial Computations
As we discussed above, the stochastic acceleration factor test is a necessary condition for the theoretical usefulness of SGD for a specific inverse problem. Here we describe another test related to parallelisability. Practically, stochastic methods for inverse problems all split the forward operator $K$ into smaller blocks $K_{i}$ such that $(Kx)_{i}=K_{i}x$. A fundamentally important question is whether $K_{i}x$ for $i=1,\dots,s$ can be computed as fast as $Kx$. Stochastic methods essentially serialise the problem which adds a computational overhead. The practical feasibility of a stochastic method then depends on both the stochastic acceleration factor and the parallelisability of the forward operator. For example, most implementations of the X-ray transform are parallelisable over its angles and the forward operator in parallel MRI is parallelisable over the coils.
Subset Sizes and Computational Overhead
The number of subsets allows us to interpolate between deterministic and high-variance random algorithms both of which have benefits and drawbacks. To illustrate this point we compare PGD with SGD. In every iteration of PGD, the most costly operations the evaluations of the gradient and the proximal operator. For imaging inverse problems often proximal operators with nonneglibable computational cost are used, e.g. the proximal operator of the total variation can be approximated via iterations [BT09, RC20] or a deep neural network applied [HLP21, TMTS24b]. In contrast, while SGD computes only a gradient on a subset, it does need to evaluate the proximal operator, too. Thus, an epoch of SGD has the same computational cost for the gradient, but has evaluated $n$-times the proximal operator which often leads to considerable overhead. Therefore, for most applications, the number of subsets $n$ needs to be chosen moderate to avoid this computational overhead.
4.2   Hyperparameter Tuning
Parameter tuning has been an important yet long-standing challenge in stochastic optimisation. Due to their random nature, most algorithmic parameters are usually much harder to be either selected a-priori or tuned on-the-fly, comparing to their deterministic counterparts. The typical algorithmic parameters which commonly need to be tuned include:
Step-Size Selection
Step-size selection is an important problem in stochastic optimisation, and has been a driving force behind the development of advanced stochastic algorithms, e.g. variance-reduced stochastic algorithms. On the one hand, stochastic algorithms inherit the step-size issues that deterministic algorithms exhibit, that is it requires knowing the Lipschitz constant, which may not be available, such as too expensive evaluate or highly spatially varying. On the other hand, step-sizes with theoretical guarantees are often too conservative. While in practice, step-sizes exceed theoretical limit often work and provide even faster performance. For example, in the first experiment of Section 5, step-size $1/(1.75nL_{\max})$ is used for SAGA under which the algorithm converges. Similarly, primal-dual based algorithms also require proper balancing of the primal and dual step-sizes to yield good performance. Overall, the step-size selection for stochastic optimisation algorithms needs more dedicated research.
Backtracking
While the line-search schemes can usually be implemented for deterministic gradient methods such as PGD or FISTA [BT09], applying these schemes for stochastic gradient methods are difficult in general both in theory and in practice. Traditional line search methods require evaluating the full objective is undesirable since it requires using all the data. Using the corresponding stochastic variants can perform well in practice but lacks theoretical guarantees [GSBR20]. Recently, a number breakthroughs have been made for on-the-fly tuning for the step-size of stochastic gradient methods applied in machine learning [DM23], but not yet fully adapted to the proximal setting which is prevalent in inverse problems.
Subset Size
The selection of the subset size is often a tricky task in practice. Although in theory the computational complexity is optimised for stochastic methods when tiny subset sizes are chosen, in practice one has to also consider the fact that most modern computational devices are highly parallel which means that the hardware would strongly favour larger subsets, leading to a trade-off. Currently there seems to be little study on providing practical guideline on selecting subset sizes in the literature.
Subset Pattern
Selecting the subset pattern is also essential for optimally applying stochastic gradient methods. As discussed in [TEGD20], the subset pattern influences the gradient Lipschitz constants on subset, and hence change the admissible step-size and convergence speed. The influence of the subset selection has also been discussed in [GDE22] for parallel MRI. The choice of subset pattern is reasonably well-understood for some inverse problems like full-angle 2D CT. However, for many other problems like limited-angle CT, 3D PET etc, this problem has not been resolved.
Restarting
Accelerated stochastic gradient methods do not automatically adapt to the strong-convexity or hidden restricted strong-convexity which is prevalent when using sparsity-inducing regularizers. Without restarting, these methods would still convergence fast but only initially. When approaching high-accuracy, the algorithms start to oscillate around the solution and which slows down convergence. To address these, adaptive restart schemes have been developed to accelerate momentum-based methods [TGBD18, FQ20].
Warm-starting
In practice, variance-reduced methods such as SVRG and SAGA, are slower than vanilla SGD at the start. Replacing the initial iterations of variance-reduced methods with a warm-start phase running vanilla SGD, a faster convergence at the beginning of iterations has been observed [TAK+23, KR17]. However, it is unclear how long this phase should be a-priori.
4.3   Memory Considerations
The stochastic algorithms presented here have very different memory footprints which can play a significant role in the way they are used in applications. In the following we will discuss the impact of memory requirements, and some mitigation strategies for SGD, SAG(A), SVRG, and SPDHG, though similar considerations can be applied to other algorithms.
Vanilla SGD does not require storing any additional pieces of information, and thus has the same memory footprint as GD, $\mathcal{O}(|\mathcal{X}|)$ required to store the current iterate.
SAG and SAGA, in their standard form, require storing the table of computed gradients. The sum of the elements in the table then either needs to be stored or computed anew in each iteration. Thus, the memory overhead of the standard implementation of SAG and SAGA is $\mathcal{O}(n|\mathcal{X}|)$, which can be prohibitively large if $n$ and $|\mathcal{X}|$ are large. In particular, in large scale CT the dimensionality $|\mathcal{X}|$ can be very large. However, memory requirements can be significantly lowered in certain mitigating scenarios. Consider linear models [DBLJ14, GSBR20], i.e. when $h_{i}=\overline{h}_{i}\circ A_{i}$ for all $i=1,\ldots,n$ and some linear operators $A_{i}$. This is for example the case for the least-squares data fidelity or the Kullback–Leibler divergence (see Section 1). The gradient of $h_{i}$ then satisfies $\nabla h_{i}(x)=A_{i}^{\ast}\nabla\overline{h}_{i}(A_{i}x)$. In other words, instead of storing $\nabla h_{i}(x)$ we can store $\nabla\overline{h}_{i}(A_{i}x)$. As the original SAGA, each iteration the application of $A_{i}$ and $A_{i}^{*}$, so no extra computational cost is introduced. The modified SAGA algorithm is written in Algorithm 12
Input: $x^{(0)}\in\mathcal{X}$, $\tau>0$
1 Initialise $y^{(0)}_{i}=\nabla\overline{h}_{i}(A_{i}x^{(0)})$ and $\gamma^{(0)}=\sum_{i}A_{i}^{\ast}y^{(0)}_{i}$
2 for $k=0,1,\ldots$ do
3       Sample $i_{k}\in\{1,\ldots,n\}$ uniformly at random
4       $\overline{y}=\nabla\overline{h}_{i_{k}}(A_{i_{k}}x^{(k)})$
5       ~$\tilde{\nabla}=A_{i_{k}}^{\ast}(\overline{y}-y_{i_{k}}^{(k)}\big{)}$ and $\gamma^{(k+1)}=\gamma^{(k)}+\tilde{\nabla}$
6       ~$\tilde{\nabla}h(x^{(k)})=n\tilde{\nabla}+\gamma^{(k)}$
7       $x^{(k+1)}=\operatorname{prox}_{\tau g}\big{(}x^{(k)}-\tau\tilde{\nabla}h(x^{(k)})\big{)}$
8       $y_{j}^{(k+1)}=\begin{cases}\overline{y}&\textrm{if~{}~{}}j=i_{k},\\ y_{j}^{(k)}&\textrm{otherwise}\end{cases}$
9 end for
Algorithm 12 Modified SAGA
The storage requirements of the modified SAGA are $\mathcal{O}(|\mathcal{X}|+|\mathcal{Y}|)$. Note that the sum of the gradients $\boldsymbol{\gamma}^{(k)}$ is of size $|\mathcal{X}|$, whereas individual gradients $y_{i}^{(k)}$ are of size $|\mathcal{Y}_{i}|$ and so the table of all $y_{i}^{(k)}$ can be stored with memory $|\mathcal{Y}|$. Moreover, we can exploit the linearity of the model so that as in the original SAGA only one adjoint evaluation is needed: to the difference between the new gradient and the stored gradient.
The choice of the preferred variant of SAGA depends on the imaging modality and the number of subsets chosen. For example, in CT images and data are often of similar sizes so the modified SAGA algorithm would be preferable already for small $n$. For PET, the image volumes are often a magnitude small than the data, so for the setting as in Section 1, the modified algorithm starts to become preferable for subset numbers exceeding 15.
The standard implementation of SVRG (see Algorithm 9) requires storing the reference iterate $\overline{x}^{(j)}$ and the full gradient $\overline{\gamma}^{(j)}$, resulting in a memory of $\mathcal{O}(|\mathcal{X}|$) which is independent of the number of subsets and comparable to GD. However, this comes at the cost of computing two stochastic gradients per iteration, one at the previous iterate and another at the reference iterate. In some applications, the memory overhead of storing $\nabla h_{i}(\overline{x}^{(j)})$ might be preferable over the extra gradient computation $\nabla h_{i}(\overline{x}^{(j)})$. As for SAG/SAGA, we can derive a modified version of SVRG which has the same memory footprint $\mathcal{O}(|\mathcal{X}|+|\mathcal{Y}|)$ and the same computational cost as the modified SAGA algorithm.
SPDHG has essentially the same memory footprint as PDHG: $\mathcal{O}(|\mathcal{X}|+|\mathcal{Y}|)$. Depending on the dual space, the memory footprint for primal-dual methods can be significantly larger compared to GD. Note that the computational effort and memory footprint is the same as the modified version of SAGA.
5   Numerical Examples
In this section we present a brief numerical study, showcasing representative stochastic optimisation algorithms discussed in this review, and comparing them against their deterministic counterparts. Note that this numerical study is not intended to provide a comprehensive and thorough comparison of algorithmic performance. Instead, our aim is to show that stochastic algorithms can indeed be beneficial for certain problems and illustrate their behaviour. The code to reproduce all results in this paper will be made available upon acceptance of this manuscript.
5.1   Sparse Spikes Deblurring
In this first experiment, we consider the toy problem of deblurringw where the ground truth are sparse spikes and compare PGD and SAGA. The purpose of this example is to find possible factors affect the performances of the two algorithms.
Let $u^{\dagger}\in\mathbb{R}^{d},d=200$ be a sparse vector, $K\in\mathbb{R}^{d\times d}$ be the matrix representation of a blur with uniform kernel of size $\kappa$ normalised such that $\|K\|=1$ and periodic boundary conditions. We consider noisy data $v=Ku+\eta$, where $\eta\in\mathbb{R}^{d}$ is random Gaussian noise. To recover $u^{\dagger}$ from the observation $v$, we minimise $\Phi(x)=g(x)+h(x)$ with $x=u$, $g(x)=\mu{|\kern-1.125pt|}x{|\kern-1.125pt|}_{1}$ and $h(x)=\frac{1}{2}{|\kern-1.125pt|}Kx-v{|\kern-1.125pt|}_{2}^{2}$. Written into the finite-sum form, we have $h(x)=\sum_{i=1}^{n}h_{i}(x)$ where $n=d$, $h_{i}(x):=\frac{1}{2}(K_{i}x-v_{i})^{2}$ and $K_{i}\in\mathbb{R}^{1\times n},i=1,...,n$ is the $i$-th row of $K$. Note that $L=1$, ${|\kern-1.125pt|}K_{i}{|\kern-1.125pt|}=1/\sqrt{\kappa},i=1,...,n$ and thus $L_{\max}=1/\kappa$.
We consider $100$ kernel sizes with $\kappa=2i-1$ for $i=1,\ldots,100$, and random Gaussian noise $\eta\sim\mathcal{N}(0,\sigma^{2}I)$, with $10$ different noise levels, taken at equidistant points in $[10^{-4},10^{-1}]$. The regularisation parameter $\mu$ is chosen as $\mu={|\kern-1.125pt|}K^{*}v{|\kern-1.125pt|}_{\infty}/3$. For PGD and SAGA, two step-sizes are chosen for each method: 1) for PGD, we choose $\tau=1/L$ (denoted PGD1) and $\tau=1.9/L$ (denoted PGD2); 2) for SAGA, we choose $\tau=1/(3nL_{\max})$ (denoted SAGA1) and $\tau=1/(1.75nL_{\max})$ (denoted SAGA2).
Given a kernel size and noise levels we first compute a high-precision solution $x^{\ast}$ and then run PGD and SAGA using the stopping criterion ${{|\kern-1.125pt|}x^{(k)}-x^{\ast}{|\kern-1.125pt|}}/{{|\kern-1.125pt|}x^{\ast}{|\kern-1.125pt|}}\leq 10^{-4}.$ We record the number of data passes needed to reach that accuracy, denoted as $N_{\text{PGD}}$ and $N_{\text{SAGA}}$ for PGD and SAGA, respectively, and compute the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ averaged over 10 noise levels. Since for each method, we have two different choices of step-sizes, therefore in total we have $4$ ratios, which are provided in Figure 1, accompanied with an extra line $1001/\kappa_{i},i=1,\ldots,$ 100.
From Figure 1, we obtain the following observations. First, when the kernel size is one, then $K$ is the identity matrix. For this case, PGD with $\tau=1/L$ converges in one step, hence the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ is very large. Also, SAGA1 is faster than SAGA2. One reason for this when for this mall kernel size, the gradients of $h_{i}$ are least correlated, and larger step-size leads to larger variance and slower performance. Second, for kernel sizes between $5$ and $199$, we observe that the ratio $N_{\text{SAGA}}/N_{\text{PGD}}$ is proportional to $1/\kappa$. Moreover, for both PGD and SAGA, larger step-size leads to faster performance. To see this, notice that the orange line $N_{\text{SAGA1}}/N_{\text{PGD2}}$ is above the red line $N_{\text{SAGA1}}/N_{\text{PGD1}}$, indicating that PGD1 needs more data pass to reach stopping criterion. While the green line below the red line means SAGA1 requires more data pass to reach the stopping criterion than SAGA2.
Figure 1: The coloured lines represent the ratio of the number of data passes for SAGA and PGD, i.e. $N_{\text{SAGA}}/N_{\text{PGD}}$, for different step-size choices. See text for more details.
5.2   Parallel CT
In this set of experiments we investigate tomographic reconstruction of a Shepp–Logan phantom of size $128\times 128$. The measurements follow the forward problem $v=Ku^{\dagger}+\eta$, where $\eta$ is a random sample from $\mathcal{N}(0,\sigma^{2}I)$, and the forward operator $K$ is defined by parallel CT with $240$ angles from the $[0,\pi)$ angle range.
We compare the performance of a deterministic algorithm PGD (Algorithm 1) with SGD (Algorithm 7) and SAGA (Algorithm 8). As discussed in Section 3, when solving inverse problems it is not common to utilise fully stochastic approaches, in the sense that it is not common to use randomly constructed batches but rather to pre-partition measurement data into batches, and then one element of the partition is sampled at random in each iteration.
PGD minimises $\Phi(x)=g(x)+h(x)$, where $x=u$, $g(x)=\lambda\text{TV}(x)$, and $h(x)=\frac{1}{2}\|Kx-v\|^{2}$. SGD and SAGA minimise $\Phi(x)=g(x)+\sum_{i=1}^{n}h_{i}(x)$, where $n\in\mathbb{N}$, and $h_{i}(x)=\frac{1}{2}\|K_{i}x-v_{i}\|^{2}$ is selected such that $\sum_{i=1}^{n}h_{i}(x)=h(x)$. This means, we define $K_{i}$, for $1\leq i\leq n$, as the CT operator that uses $240/n$ angles from the range $[i/n,\pi)$ by taking every $n$-th angle starting with $i/n$ giving a $43n/$ 4-degree angle separation. This is sometimes called the equidistant or staggered partition, and results in all the operator norms $\|K_{i}\|$ to be of a nearly identical value. Moreover, the number of batches is typically a divisor of the number of angles, which ensures that all the batches are of the same size.
(a) Objective value over 200 data passes
(b) Objective value over the first 20 data passes
(c) Reconstructions after 10 data passes with $n=60$
Figure 2: Performance comparison for a Shepp–Logan phantom from CT measurements corrupted by Gaussian noise with $\sigma=1$. Note that the objective value for SAGA starts at the $1$-st data pass due to initialising the table of gradients at $x^{0}$.
The initial image is $x^{0}=0$. The step-size for PGD is $\tau=1/\|K\|^{2}$. For SGD we use a decaying step-size schedule, $\tau_{k}=1/(2nL_{\max}(1+0.01k/n))$, where $L_{\max}=\max_{i}\|K_{i}\|^{2}$, following [TAK+23], which satisfies condition (25). In this regime the step-sizes decay from $1/(2nL_{\max})$ to roughly $1/(4nL_{\max})$, over the studied $100$ data passes, giving a behaviour similar to a constant step-size regime, albeit convergent. For SAGA we use $\tau=1/(3nL_{\max})$ as step-size.
In Figure 2 we show the performance of these 3 algorithms in terms of the sub-optimality of the objective value: $\Phi(x^{(k)})-\Phi(x^{\ast})$, where $\Phi(x^{\ast})$ is a reference solution computed with FISTA. Figure 2 shows a significant speed up with SGD and SAGA over PGD. SGD is the fastest initially, but it then slows down after 5-10 data passes. On the other hand, SAGA is steady and fast throughout the optimisation trajectory and the increased variance it presents in initial iterations disappears already after around 20 data passes. Figure 2a and Figure 2b also investigate the effect of the number of subsets: the lightest shade shows a lower number of subsets ($n=10$), normal shade shows ($n=60$), and the darkest shade shows a high number of subsets ($n=240$). We observe that increasing the number of subsets has a dual effect: it increases the variance of the optimisation trajectory for stochastic algorithms, but it greatly increases the speed. In Figure 2c we show the reconstructions of the 4 studied algorithms obtained after $10$ data passes. The reconstruction with PGD presents as oversmooth, and fails to adequately reconstruct the finer image detail. For example, consider the zoomed in region where stochastic algorithms manage to recover two distinct objects (one of a lower and the other of a higher intensity), whereas PGD does not.
(a) SGD with $n=30$
(b) SAGA with $n=30$
(c) SAGA with $n=240$
Figure 3: Comparison of reconstruction performance for a walnut from fan-beam CT measurements with three noise levels
In Figure 3 we compare the effect of the strategy for accessing the function indices. Specifically, we compare SGD and SAGA, with indices sampled uniformly at random (with replacement), and IGD and IAGA, which are incremental gradient modifications of SGD and IAGA that access indices according to a deterministic Herman–Meyer order [HM93]. The results show that the behaviour highly depends on the number $n$ of subsets and the algorithm. Comparing non-variance reduced algorithm reveals that the deterministic strategy can outperform the stochastic strategy, though the differences are not significant. However, for variance reduced algorithms the differences are significantly more severe. When $n$ is small, SAGA outperforms IAGA and presents a much smaller variance, though the variance with IAGA reduces significantly within the first $50$ data passes. When $n$ is large, SAGA is significantly better than IAGA.
5.3   Fan-Beam CT
In the last set of experiments we take the central slice of a cone beam CT scan of a walnut [Mea20]. Full sinogram data has 721 projections in the $[0,2\pi)$ range. To reduce the computational costs we consider only $360$ projections in the $[0,2\pi)$ range, with a $1^{\circ}$ angle separation. The reconstructed images are of size $280\times 280$.
Noisy measurements are generated by corrupting the sinogram (the measurement data) following the Beer–Lamberts law. We first compute the expected number of counts via $N_{I}=\text{Pois}(I_{0}\exp(-v))$, where $v$ is the loaded (noiseless) sinogram and $I_{0}$ the intensity of the initial beam. The noisy measurement is then computed by first correcting all 00 entries in $N_{I}$ to $1$, and then applying linearisation (post-log measurements) as $v^{\delta}=-\ln(N_{I}/I_{0})$. We consider three intensity levels: low intensity ($I_{0}=50$), medium intensity ($I_{9}=250$), and high intensity ($I_{0}=5000$), simulating high to low noise. The objective is defined as
$\Phi(x)=\frac{1}{2}\|Kx-v\|_{2}^{2}+\alpha_{I}\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x),$ (29)
where $\alpha_{I}\in\{0.124,0.045,0.01\}$ is the regularisation parameter for high, medium, and low noise levels, respectively. Reference solutions are shown in Figure 4.
Figure 4: Reference solutions for high, medium and low noise levels
We study gradient based methods (PGD, SGD, SAGA, SVRG) and primal-dual methods (PDHG and SPDHG). The terms in the objective (29) are identified with optimisation templates (7) and (22), for gradient based and primal-dual methods as follows
Gradient based methods
$g(x)=\alpha_{I}\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $f\equiv 0$, and $h(x)=\frac{1}{2}\|Kx-v\|^{2}$ for PGD and $h_{i}(x)=\frac{1}{2}\|K_{i}x-v_{i}\|^{2}$ for SGD, SAGA and SVRG;
Primal-dual methods
$g(x)=\alpha_{I}\|\nabla x\|_{1}+\iota_{[0,\infty)^{d}}(x)$, $A=K$ with $f(x)=\frac{1}{2}\|x-v\|^{2}$ for PDHG and $A_{i}=K_{i}$ with $f_{i}(x)=\frac{1}{2}\|x-v_{i}\|^{2}$ for SPDHG, $h\equiv 0$.
The proximal operator of $g$ is approximated by 100 FGP iterations [BT09]. Stochastic methods use $n=60$ (for SGD, SAGA and SVRG), and $\ell=60$ (for SPDHG), subsets of the data. Operators $K_{i}$ are defined analogously to the parallel CT case in Section 5.2: starting with the angle $i/n$ (or $i/\ell$) we take every $n$ (or $\ell$) angle from the equidistant partition of $[0,2\pi)$ into $360$ angles. As a result, the norms $\|K_{i}\|$ are all nearly the same. SGD uses a constant step-size $\tau=1/(2nL_{\max})$, and SAGA and SVRG use $\tau=1/(3nL_{\max})$. SPDHG is implemented following [EMS19, Algorithm 2], with $\sigma=\gamma\rho/K_{\text{max}}$, $\tau=1/(n\gamma K_{\text{max}})$, with $\rho=0.99$, $\gamma=1$ and $K_{\text{max}}=\max_{i}\|K_{i}\|$.
The results in Figure 5 show that the noise level can have a noticeable effect on the convergence behaviour. First, for lower noise levels SGD is competitive with PGD within the given optimisation window, whereas for higher levels it is less competitive. On the other hand, SAGA and SVRG are clearly ahead of PGD regardless of the noise level. Second, SPDHG is the fastest method for lower noise, but its performance decays (relative to SVRG and SAGA) as the noise increases.
(a) High noise
(b) Medium noise
(c) Low noise
Figure 5: Comparison of reconstruction performance for a walnut from fan-beam CT measurements with three noise levels
In spite of the superior convergence behaviour of SAGA, SVRG and SPDHG, running a large number of iterations might not be desirable due to computational or time constraints or even needed if the sought features of the image are recovered early. It is therefore of interest to investigate which algorithms recover a good quality reconstruction within a limited iteration budget, which image details are missing, and whether an increased computational or memory budget is justified or whether SGD would suffice. To investigate this, we look at reconstructions with SGD, SVRG, SAGA, and SPDHG within the first $10$ data passes.
The results in Figure 6 confirm that all the stochastic algorithms converge and recover a visually satisfying solution quickly. Moreover, the variance reduced algorithms shows a steady convergence behaviour and better quality of detail. However, SGD also recovers a good quality solution, and depending on the problem, only a handful of data passes might be needed to recover general features of the solution. Thus, a sound strategy to improve the performance of an image reconstruction method would be to warm start the reconstruction with a few data passes with standard SGD, before employing a more advanced method. This would also help reduce the initially high variance seen with SAGA, which updates a gradient stored in the table only once the corresponding batch (or subset) index is selected (which introduces an initial bias towards selected subset gradients) and may result in a reconstruction of unreliable quality within the first $10$ data passes.
As argued in Section 4, the step-size regimes are often pessimistic, in the sense that they are intended to serve as the largest value that ensures convergence towards a minimum, with a given method, for all admissible images. In Figure 7 we investigate this behaviour by looking at the convergence trajectory of SVRG for medium noise ($I=250$), with $n\in\{30,60,120\}$ and constant step-sizes $\tau=\delta/(nL_{\max})$ for $\delta\in\{1/3,1,2\}$. The last step-size was chosen such that it is the largest for which SVRG still converged to a good solution. All three batch sizes stopped converging for $\delta=7/3$. The results show that SVRG is stable with respect to the step-size choice. It gets progressively more unstable as the number of subsets $n$ increases: when $n=120$, SVRG performs best for $\delta=1$ and the worst for $\delta=2$, whereas for $n=30$ it performs the best for $\delta=2$ and the worst for $\delta=1/3$. This behaviour is (roughly) consistent across the studied noise levels. Note that the same experiment was carried out for SAGA, which showed divergent behaviour already for $\delta=1$.
(a) SGD
(b) SVRG
(c) SAGA
(d) SPDHG
Figure 6: Reconstruction with SGD, SVRG, SAGA, and SPDHG for medium noise. The colour scheme in all the images is such that the smallest and largest value corresponds to the maximum and minimum values in the reference solution.
Figure 7: Behaviour of SVRG with for the reconstruction of the walnut with respect to the step-size and number of subsets $n$.
6   Outlook
Optimisation algorithms are a backbone for solving inverse problems with stochastic approaches becoming increasingly relevant for large-scale applications. We have discussed state-of-the-art algorithms for inverse problems with a particular focus on the toolbox nature of the topic. Virtually all relevant algorithms are assembled from a fairly small range of tools from both the deterministic and stochastic optimisation toolkit. Particularly the stochastic toolkit has been developed with machine learning in mind but it turns out to be of excellent use for inverse problems, too.
Over the next decade, we expect this trend to continue, with more and more applications realising the potential that stochastic optimisation offers. Furthermore, we expect stochastic optimisation tools to be developed specifically for inverse problems which may not be relevant to machine learning. Potential examples are sketching in the image domain and nonconvexity from nonlinear forward operators. We also expect that theory could be developed specifically for inverse problems. For instance, it has been shown that stochastic gradient descent can be much faster in the so called interpolation regime [VML+19]. In the context of inverse problems this regime implies that there is no noise and no regularisation which of course is unrealistic. Instead, there could be theory tailored to inverse problems explaining which algorithms are fastest in a specific noise regime, thereby potentially explaining the observations we made in Section 5.
Finally, as we already indicated throughout the manuscript in various further reading sections, an exciting path forward is the interplay between inverse problems, optimisation and machine learning. This could be in the context of modelling like learning regularisation models or learning forward operators or their sampling, but also learning optimal algorithms to solve inverse problems. We believe that the current optimisation-based paradigm with learned explicit regularisation will become a dominating computational paradigm in inverse problems, while developing efficient optimisation algorithms tailored for these applications will continue to be important to practitioners. To the best of our knowledge this has not been attempted in the context of stochastic algorithms. In our view this development is far from over and more research is needed to fully understand and exploit the connections in the context of inverse problems.
We acknowledges support from the EPSRC: MJE (EP/Y037286/1, EP/S026045/1, EP/T026693/1, EP/V026259/1) and ZK (EP/X010740/1). JL is supported by the “Fundamental Research Funds for the Central Universities”, the National Science Foundation of China (BC4190065) and the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102).